A Complete Bibliography of Philosophy of Science: 1950–1959

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

13 March 2017 Version 1.02

Title word cross-reference

Ψ [140].

-Function [140].

1 [635]. 1900 [322]. 1950 [588]. 1954 [423].

20-21 [369].

Abetti [253, 253]. Ability [214]. Abraham [459, 464]. Abramson [235, 370]. Abstraction [15]. Abstracts [377, 527, 543, 568, 614, 613, 659, 709, 375]. Academic [717]. Acceptance [474]. Account [211]. Ackoff [317, 364]. Action [70, 353, 287]. Activity [191]. Adrenal [410]. Adventure [289]. Aesthetics [215, 407]. Affiliations [428]. after [688]. Age [633]. Ages [610]. Aggregate [167]. Agnes [408].

Agree [233]. Agreement [597]. Aid [676]. Aiken [633]. Aim [348]. Alan [593]. Albert [345]. Alchemy [371]. Alexander [719, 562]. Alfred [558, 327]. Alice [23]. Allen [258]. Alternative [342]. Ambrose [23]. America [147]. American [64, 365, 629]. among [385]. Ampère [41]. Amplifying [114, 355]. Analogy [101, 487]. Analysis [663, 412, 211, 494, 414, 178]. Analytic [665, 143]. Anatomy [651, 199]. Ancient [198]. Anderson [459, 593]. Andre [653]. Anscombe [344, 670]. Anthropology [10, 521, 576, 73, 503, 443]. Anti [536]. Anti-Matter [536]. Antinomy [691]. Antiquity [674]. Anton [366]. Anwendungen [390]. Application [243, 505, 160]. Applications [471, 372, 131]. Applied [280, 9]. Appraisal [547]. Approach [353, 243, 164, 710]. Approaches [689]. Arber [408]. Areas [238]. Argument [665, 677]. Argyle [608]. Aristotle [586, 522]. Arne [635]. Arnold [367, 714, 407]. Art [268]. Arthur [391, 656, 39, 75, 143]. Article [405]. Articles [196]. Ashley [177]. Aspects [181, 206, 411, 606, 296]. Association [684]. Astrology [371]. Astronomy [286, 626, 253]. Atomic [699, 672]. August [423]. Auguste [646]. Augustus [392]. Austin [234]. Authorship [698]. Autiobiography [173]. Autobiography [62]. Automation [541]. Available [355]. Avrum [465]. Axiomatic [281]. Axiomatization [340]. Ayer [561].

B [232, 261, 262, 290, 542, 363, 199, 260, 480, 253, 407]. Back [273, 293, 307, 334, 351, 378, 396, 416, 435, 445, 468, 485, 497, 515, 528, 545, 570, 580, 594, 616, 637, 660, 682, 696, 724]. Backwards [295]. Bacon [244, 251, 146]. Baitsell [330]. Baker [202]. Band [24]. Barker [704]. Barnes [89]. Basic [560, 46, 441]. Basis [599, 591, 517, 1]. Baumgart [195]. Bayliff [365, 365]. Be [523]. Behavior [368, 35, 87, 27, 49, 50]. Behavioral [647, 14, 53]. Behaviorism [664, 208]. Behaviour [608]. Being [177]. Belief [566, 519, 488]. Believer [715]. Bender [55]. Bendix [428]. Benjamin [542, 25, 146, 509]. Bentley [391]. Bergmann [389]. Bergson [78, 188]. Berkeley [326]. Bertalanffy [24, 552]. Bertram [611]. Bertrand [588, 36, 269]. Berücksichtigung [390]. Beryl [407]. Besonderer [390]. between [101, 206, 712]. Biochemical [591]. Biological [265, 599, 211]. Biologische [24, 602]. Biology [99]. Bjorset [60]. Black [414, 43, 178, 254]. Blaine [365]. Blair [483]. Blochman [653]. Blyth [671]. Boas [563]. Bochenski [61]. Body [154, 210]. Bohr [672]. Book [62, 63, 64, 78, 110, 144, 175, 181, 197, 198, 216, 218, 232, 233, 258, 259, 261, 262, 263, 264, 265, 266, 267, 268, 269, 270, 271, 289, 290, 291, 321, 325, 326, 327, 328, 132, 37, 345, 346, 347, 367, 368, 369, 161, 366, 695, 718, 129, 508, 511, 542, 588, 589, 363, 605, 22, 41, 180, 538, 553, 555, 509, 560, 651, 693, 65, 145, 199, 158, 182, 195, 426, 159, 23, 196, 260, 670, 201, 118, 633, 557, 414, 713, 24, 40, 646, 717, 332, 412, 413, 91, 10, 481, 482, 650, 57, 317, 657, 25, 149, 391, 322, 459, 541, 607, 612]. Book [647, 648, 681, 59, 179, 203, 219, 464, 465, 510, 609, 610, 611, 58, 656, 94, 160, 177, 235, 370, 374, 671, 677, 318, 371, 373, 375, 408, 410, 428, 429, 430, 431, 480, 591, 592, 61, 92, 131, 694, 365, 393, 411, 608, 627, 652, 389, 458, 561, 587, 676, 60, 427, 251, 55, 606, 631, 673, 714, 329, 331, 483, 539, 556, 632, 634, 711, 655, 147, 213, 257, 323, 330, 348, 372, 392, 563, 672, 130, 148, 234, 77, 200, 252, 254, 217, 590, 593, 653, 654, 26, 512, 715, 629, 390, 716]. Book [710, 109, 174, 202, 214, 146, 364, 635, 562, 692, 720, 128, 344, 406, 479, 253, 286, 564, 678, 172, 38, 674, 540, 649, 558, 176, 215, 287, 324, 394, 461, 462, 624, 27, 457, 42, 320, 39, 75, 143, 255, 554, 559, 680, 721, 319, 93, 43, 173, 460, 56, 89, 76, 409, 463, 628, 675, 679, 722, 90, 79, 407, 288, 625, 183, 712, 719, 256, 626, 9, 630, 178]. Books [95, 111, 236, 272, 292, 306, 333, 349, 376, 395, 415, 432, 466, 513, 565, 658]. Born [264, 22]. Bouwsma [407]. Boyle [452]. Brain [118, 266]. Braithwaite [363]. Brameld [710]. Bridgman [110, 213]. Bright [347]. British [527, 543, 568, 614, 631]. Bronislaw [681]. Brown [629, 291, 203, 627]. Brownell [202]. Brumbaugh [176, 394]. Brunswick [412]. Brynjolf [60]. Budget [392]. Burali [622]. Burali-Forti [622]. Burniston [203]. Business [447]. Butler [564, 234]. Butterfield [259].

C [216, 321, 327, 132, 345, 145, 413, 317, 149, 561, 251, 631, 711, 257, 372, 130, 148, 42]. Cajal [118]. Caldin [58]. Campbell [256]. Can [312, 523, 295, 233]. Cancer [494]. Cannon [118]. Cantrill [175]. Carl [219]. Carnap [141, 473, 512, 390]. Carol [131]. Carola [195]. Carroll [542]. Cartwright [525]. Case [262]. Cassirer [430]. Cassius [271]. Categories [552, 418]. Causal [265, 211]. Causation [154]. Cause [22]. Causes [193]. Cautions [701]. Cells [702, 494]. Century [679, 631]. Chall [375]. Challenge [30]. Chance [695, 22]. Change [312]. Changing [449]. Charles [588, 160, 540, 324, 93]. Chase [109]. Chaudhury [458]. Chauncey [233, 169, 722]. Chemistry [322, 600]. Chisholm [630]. Choice [695]. Christian [556]. Churchman [317]. Circle [252]. Circuits [621]. Circular [265]. Civilization [132, 694, 217]. Clarification [663]. Clark [365, 319]. Clarke [562]. Class [34]. Class-Membership [34]. Classical [74, 340]. Clinical [380]. Clock [668, 314, 359, 669, 404]. Closed [84]. Clyde [10, 628]. Co [507]. Co-Presence [507]. Code [557]. Cohen [331]. Collection [196]. Collins [462, 318]. Combinatory [550, 621]. Commandments [219]. Comment [227, 552, 537, 54]. Comments [477, 226, 276, 17, 405, 383, 526, 284, 48, 404]. Communication [493]. Community [481, 202, 76]. Comparative [383, 56, 382]. Compass [394]. Complete [140]. Composition [425]. Computers [648]. Conant [262, 232]. Concept [223, 280, 74, 549, 53, 585, 456, 591]. Conception [167, 714, 210, 48, 504]. Conceptions [600]. Concepts [278, 82, 560, 239, 619, 213, 673, 457]. Conceptual [532, 281, 508]. Concerning [282, 506, 573, 596]. Condorcet [624]. Conference [369, 611, 370]. Confirmation [640, 6, 707, 473, 705, 442, 641]. Confirmatory [403]. Conflict [361, 324, 428]. Conflicting [40]. Conger [183]. Congress [423]. Connaissance [172]. Connected [225]. Consciousness [505, 370, 413, 235]. Considerations [596]. Consistency [281]. Consistent [167]. Constructions [5]. Contemporary [102, 223, 665, 500, 393]. Content [103]. Continuity [700, 207, 558]. Continuum [167]. Contributions [437, 412]. Control [316, 421]. Conventional [319]. Cooperative [106]. Copenhagen [701]. Copernican [626]. Copi [479, 479]. Cornelius [509]. Cornu [646]. Correction [566]. Correlation [154]. Correspondence [562]. Corrigibility [277]. Cortex [410]. Cosmic [536]. Cosmogony [628]. Cosmological [574]. Cosmologies [135]. Cosmology [623, 136, 606, 179]. Coulson [711]. Council [629]. Counterparts [583]. Creation [606]. Creative [629, 181]. Crime [160]. Criminalistics [160]. Crisis [189, 717, 406]. Critic [406]. Critical [341, 692]. Criticism [704]. Critics [701]. Critique [216, 172]. Cross [597]. Cross-Cultural [597]. Cultural [597, 607, 298, 73, 503, 710]. ’Culture [248, 163, 166, 238, 681, 341, 694, 249, 129, 655]. Current [240, 139]. Currents [346]. Curry [628]. Cybernetics [265, 369].

D [233, 269, 321, 633, 647, 561, 483, 556, 217, 649]. Dagobert [269, 94, 217]. Dagut [675]. Dantzig [406]. D’Arcy [210]. Darlington [707, 705]. Darwin [384]. Datum [502]. David [373, 365, 392, 629]. Davidson [538]. DeBroglie [538]. Deciding [702]. Decision [339, 475]. Decisions [71]. Deductive [427]. Deeper [676]. Defining [473]. Definition [572, 687, 505, 442]. Definitions [507, 47, 248, 163, 249]. Degree [168, 473, 442]. Denise [345]. Denonn [269]. Denotation [718]. Descartes [41, 693]. Description [140]. Design [258, 648, 364]. Designation [691]. Detection [160]. Determinism [700]. Deutsch [629]. Develop [214]. Development [263, 221, 626, 291, 592]. Dewey [25, 391]. Dialectic [400, 309, 337]. Dialectica [613, 659, 709]. Dialectical [450, 618]. Dialetical [501]. Dictionary [269, 409, 269]. Difficulties [46]. Digital [648]. Dimension [14]. Dingle [289]. Direction [693]. Discipline [387]. Discourse [358, 510]. Discovery [212]. Discussion [501, 156]. Disease [310, 720]. Disposition [53]. Dispositional [398]. Do [357]. Dobzhansky [332]. Dominant [563]. Donald [589, 304]. Dorothy [118]. Douglas [366]. Dr [173]. Dr. [477, 360, 405, 525, 249]. Dryer [286]. Dubarle [556]. Duhem [348]. Duhemian [665]. Duncan [234]. Duncan-Jones [234]. Dynamical [440]. Dynamics [673].

E. [537]. East [31]. Easton [365, 251]. Ecology [27]. Economic [551]. Economics [437]. Economist [366]. Eddington [39, 79]. Edel [464]. Edgar [75, 276]. Edition [372]. Editor [478]. Edmund [590]. Education [710]. Edward [553, 722]. Edwards [258, 510]. Effects [193]. Efficacy [357]. Effort [27]. Einführung [390]. Elaine [410]. Elementary [453]. Elements [167, 143]. Elmer [89]. Else [412]. Elton [407]. Emergence [541]. Emil [173]. Emotion [713]. Emotive [465]. Empirical [642, 244, 103, 517]. Empiricism [13, 263]. Empiricist [719]. End [597]. Energy [355]. Engineering [182]. England [430, 679]. Entities [5]. Entropy [315]. Environments [531]. Epicurus [463]. Epistemological [505, 547, 504, 1].

Epitomization [183]. Equivalence [125]. Eric [715]. Erich [413]. Erik [713]. Erkenntnis [602]. Ernest [587, 625]. Ernst [430]. Essay [676, 331, 418]. Essays [588, 634, 625, 657, 391, 9]. Essence [523]. Essentials [144]. Establishment [506, 472]. Esthetics [599]. Ethel [345]. Ethical [276, 597, 719, 464, 539]. Ethics [599, 448, 32, 488, 619, 345, 612, 464, 465, 719]. Étude [172]. Euclid [79, 586]. Eugene [365, 392]. Europäische [61]. Evaluating [291]. Evaluation [681, 31]. Events [230]. Evidence [578, 403]. Evil [483]. Evolution [135, 639, 411, 57, 610]. Exact [137, 674]. Examination [381, 470]. Examined [470]. Existential [3, 20]. Existentialism [115]. Existenzialismus [38]. Experience [662, 552, 175, 25]. Experiment [662, 105]. Experimental [262, 69, 391, 88, 258, 657]. Experimentation [316]. Explanation [363, 279, 520, 521, 603, 440]. Explanations [438]. Explanatory [357]. Explicandum [74]. Explorer [118]. Extended [167, 604]. Extensive [15]. Extra [699]. Extra-Atomic [699]. Eye [408].

F [327, 65, 118, 717, 413, 391, 58, 561, 539, 130, 629]. Fact [582]. Factor [316]. Factors [492, 526]. Factual [168]. Falsity [685]. Farber [37, 201]. Farrington [146]. Feedback [265]. Feigl [20]. Ferm [161]. Ferrier [41]. Figures [170]. Finlay [179]. Finlay-Freundlich [179]. Firth [681]. Fisher [181]. Five [13, 701]. Five-Fold [13]. Flew [196]. Flexible [618]. Flow [295]. Foerster [265, 369]. Fold [13]. Forbes [182]. Force [673, 139]. Form [103]. Formal [386, 387]. Formalism [704, 534]. Formalization [35, 585]. Forti [622]. Fothergill [411]. Foundations [673, 548, 710, 554, 460, 719, 558]. Founders [57]. Four [170]. France [201]. Francis [593, 244, 146]. Frank [716, 409, 712, 643, 77]. Fraser [647]. Frederick [198, 325]. Free [530, 64]. Free-Will [530]. Freedom [354, 666]. Freeman [431]. Frege [254]. Frenkel [412]. Frenkel-Brunswick [412]. Freundlich [179]. Friedmann [541]. Friendly [450]. Front [11, 29, 45, 68, 80, 97, 112, 120, 133, 150, 162, 186, 204, 220, 237, 275, 294, 308, 335, 352, 379, 397, 417, 436, 446, 469, 486, 498, 516, 529, 546, 571, 581, 595, 617, 638, 661, 683, 697, 725]. Function [363, 140]. Functions [687, 138]. Fundamental [443]. Fundamentalism [618]. Fundamentals [23]. Further [501]. Furthering [299]. Future [52, 37].

G [695, 65, 196, 670, 203, 610, 131, 411, 561, 330, 372, 653, 562, 344, 564, 461]. Gallie [407]. Games [424, 311]. Gardner [426]. Garnett [256]. Gaston [624]. Gaynor [716, 409]. Geach [254]. Gegenwart [61]. Geiringer [721]. General [437]. Genetics [332]. Genetotrophic [591]. Georg [428, 680]. George [267, 539, 563, 27, 183, 419]. Georges [541]. Gerhart [290]. Germany [503]. Gestalt [155, 246, 245, 487]. Gibson [93, 218]. Gilbert [407]. Gilles [624]. Gilles-Gaston [624]. Ginsberg [90]. Giorgio [253]. God [711]. Goethe [158]. Good [483]. Goodman [456]. Götlind [713]. Gotshalk [649]. Gottlob [254]. Goudge [149]. Government [64]. Gramlich [593]. Granger [624]. Graubard [371]. Gravitation [180].

Grazia [76]. Great [345]. Greeks [675]. Grimes [365]. Gross [327]. Group [611, 428, 611]. Group-Affiliations [428]. Grünbaum’s [404]. Guide [267]. Guilt [290]. Gustav [389, 38]. Guterman [612].

H [78, 265, 369, 693, 65, 145, 322, 235, 370, 428, 92, 365, 587, 147, 419, 562, 722, 719]. Hadley [175]. Hammond [101]. Hampshire [407]. Handbook [426]. Handy [526]. Harcourt [629]. Hardt [219]. Harold [693, 541, 559]. Harris [693]. Harry [89]. Harvard [262]. Hawkins [629]. Hebrew [217]. Heinrich [38]. Heinz [158]. Heisenberg [714, 165, 323]. Helen [650, 407]. Henle [170]. Henri [406]. Henrietta [716]. Henrik [680]. Henry [605, 260, 633, 671, 131, 372, 320]. Herbert [259, 289]. Herdan [695]. Hesse [480]. Heuristic [83]. Hilda [721]. Hildebrand [65]. Hindsight [403]. Historical [152, 411]. Histories [262]. History [198, 267, 346, 161, 4, 182, 438, 413, 520, 575, 676, 147, 130, 716, 253, 286, 666, 457, 89, 576, 563, 462, 197]. Hochberg [477]. Hoffer [715]. Hofmann [716]. Holland [107]. Hope [91]. Horace [320]. Hourani [539]. Hugues [427]. Hull [413, 35]. Human [511, 118, 481, 541, 373, 480, 655, 672, 202, 540, 27, 639, 177, 27]. Humanism [556]. Hume [321, 325]. Hunter [215]. Hutten [587]. Hypotheses [3, 474, 20]. Hypothesis [641].

Idea [65, 711]. Ideal [533, 84]. Ideas [535, 505, 147, 299, 676]. Identity [650]. Ideological [324]. Ideology [633]. Idiom [577]. Ihrer [390]. II [337]. Imagination [480]. Impact [217]. Imperatives [685]. Impertinent [4]. Implications [620]. Importance [500]. Incongruity [583]. Indeterminacy [165, 354]. Indeterminism [690]. Indications [385]. Indirect [358]. Individualism [362]. Individualisms [572]. Individuality [591]. Induction [222, 36, 51, 680]. Inductive [315]. Industrial [296, 146, 541]. Inference [524, 88, 559]. Inferred [5]. Information [12, 67, 81, 119, 134, 185, 205, 274, 350, 434, 496, 569, 636, 723, 381, 315, 439]. Inquiries [391]. Inquiring [491]. Inquiry [30, 317, 391, 508, 609, 635]. Insight [380]. Institution [72]. Integrated [84]. Intellectual [492, 526]. Intention [670]. Intentional [260]. Interdisciplinary [710, 635]. International [423]. Interpretation [520, 450, 701, 604]. Interpretations [3, 17]. Interpretative [596]. Interrelational [649]. Interrogatives [685]. Introduction [347, 317, 160, 671, 587, 427, 215, 27, 89, 288, 555, 479, 56]. Intuitionism [534]. Investigations [344]. Irving [479]. Isenberg [407]. Isomorphisms [126]. Issue [139]. Ivan [657].

J [218, 271, 37, 605, 65, 182, 591, 561, 714, 720, 286, 564, 172, 554, 460, 407]. Jacks [39]. Jacques [172, 56]. Jakob [180]. James [218, 232, 262, 160, 318, 430, 393, 242]. James-Lange [242]. Jammer [673, 457]. Jarrett [393]. Jasper [38]. Jaspers [59]. Jean [54]. Jeffreys [559]. Jennings [365]. Jessop [326]. Jivan [458]. Joachim [693]. Joad [216]. Joergen [263]. Joergensen [263]. Johannes [195]. John [270, 542, 557, 25, 391, 671, 329, 257, 174]. Johnson [589]. Jones [234]. Jørgensen [219]. Joseph [716, 319]. Journal [527, 543, 568, 614, 635]. Judgement [483]. Judges [380]. Judgment [464, 462, 589]. Judgments [298, 361, 187]. Jurisprudence [69]. Justification [36].

Kantianism [2]. Karl [560, 40, 59, 629, 172, 38]. Kathleen [431]. Kattsoff [712]. Kemble [227]. Kenneth [214]. Kepler [195, 286]. Kershner [199]. Kester [628]. Key [207, 139]. Keyes [214]. Keyser [271]. Kierkegaard [318]. Kinship [533]. Klickstein [322]. Kluckhohn [10]. Kneale [561]. Knight [407]. Knowledge [588, 105, 122, 676, 483, 552, 692, 321, 412, 672]. Kogan [144]. Koren [605]. Körner [508]. Kraft [252]. Kuhn [626]. Kurt [428].

L [258, 199, 317, 541, 629, 364, 286, 39]. L. [552]. Laboratory [598]. Lachman [554]. Ladd [557]. Laer [605]. Lake [407]. Lamont [483]. Landé [558]. Lange [242]. Langford [145]. Language [502, 238, 196, 206, 533, 587, 632, 311, 564, 255, 407, 695, 542, 43]. Languages [442]. Larrabee [78]. Latency [604]. Lavoisier [366]. Law [363, 211, 277, 181]. Lawrence [653, 592]. Laws [279, 640, 312, 357, 551, 707, 705]. Layman [140]. Lazare [634]. Lazarsfeld [717]. Lazerowitz [23]. Leaf [360]. Leake [233]. Learned [629]. Learning [341]. Learning-Theory [341]. Least [27]. Lebens [24]. Leblanc [708, 427]. Lee [542]. Leibniz [562, 470, 429]. Leibniz-Clarke [562]. Leicester [322]. Lelland [720]. Leo [375]. Leonard [612, 365, 255, 460, 671]. Lepley [632]. Leslie [129, 694]. Lester [269]. Letter [478]. Letters [195, 94]. Lewis [268, 145]. Liberty [200]. Lies [685]. Life [165, 195, 118, 26, 720, 649, 10, 55, 42, 75, 590]. Light [489]. Limits [58, 33]. Lindley [249]. Lindsay [373]. Lindzey [426]. Line [138]. Linear [167]. Linguistic [225, 703, 126, 107]. Linguistics [409]. Linhart [706]. Link [712]. Lins [609]. Linsky [255]. Little [481]. Living [114, 210, 320]. Loehrich [374]. Logic [82, 300, 550, 621, 521, 510, 200, 169, 315, 439, 319, 603, 440, 479, 145, 23, 260, 391, 671, 427, 512, 128, 479, 588, 196, 320, 625]. Logical [216, 263, 508, 13, 115, 16, 21, 30, 242, 389, 680, 456, 648]. Logik [390]. Lossky [197]. Louis [538, 712]. Loyd [365]. Ludwig [24, 344]. Lydia [429]. Lynd [650].

M [63, 216, 37, 345, 367, 718, 589, 670, 322, 61, 344, 479, 630]. Mace [631]. Mach [667]. Machine [225]. Macklem [651]. MacPartland [270]. Madden [245, 722]. Magnus [158]. Main [346]. Maker [182]. Makes [187]. Making [64, 475]. Malinowski [681]. Malinowskian [521]. Man [175, 232, 10, 694, 653, 720, 39, 256, 182, 681, 60]. Mandelbaum [593]. Mandelker [180]. Manheim [59]. Mankind [109]. Mannheim [172]. Maquet [172]. March [270, 369]. Margaret [407]. Margenau [131, 372, 2, 98]. Marie [462]. Mario [609, 431, 409]. Marjorie [91]. Mark [371]. Marquette [56]. Marquis [624]. Marsh [588]. Martin [538, 718]. Marvin [201, 655]. Marxian [646]. Mary [480]. Mason [327, 64, 346]. Mass [715]. Materialism [450, 501]. Mathematical [281, 689, 288]. Mathematics [387, 471, 553, 560, 199, 634, 716, 461]. Mathématique [624]. Matrix [224]. Matter [269, 11, 29, 45, 68, 80, 97, 112, 120, 133, 150, 162, 186, 273, 204, 220, 237, 275, 293, 307, 334, 351, 294, 308, 335, 352, 378, 396, 416, 435, 379, 397, 417, 436, 445, 468, 485, 497, 446, 469, 486, 498, 515, 528, 545, 570, 516, 529, 546, 571, 580, 594, 616, 637, 581, 595, 617, 638, 660, 682, 696, 724, 661, 683, 697, 725, 114, 536, 241, 270, 257]. Maurice [593]. Max [62, 264, 22, 414, 673, 254, 457, 42, 43, 178]. Maxwell [440]. Mayer [198]. Maziarz [553]. McDonald [407]. McGill [37]. McKie [366]. McKinley [610]. McKinney [228]. McMurrin [393]. Mead [419, 215]. Meaning [338, 190, 331, 573, 14, 116, 512]. Means [597]. Measurement [113, 573, 549, 604]. Measuring [138]. Mechanical [604]. Mechanics [114, 2, 113, 207, 340]. Mechanism [104]. Mechanisms [265]. Mechanistic [699, 48]. Medicine [137]. Medieval [198, 86]. Meeting [31]. Members [684]. Membership [34]. Memory [188]. Men [312]. Mendeleev [600]. Menger [560]. Merrell [650]. Merton [675, 652]. Metaphysic [452]. Metaphysical [224, 399]. Metaphysics [389, 634, 625]. Method [367, 238, 203, 280, 15, 296, 136, 383, 382, 73, 620, 713, 317, 331]. Methodological [586, 572, 46, 362]. Methodology [686, 642, 211, 368]. Methods [647, 655, 299, 313, 317, 128]. Meyer [173, 173]. Michael [651, 608, 200, 692]. Microphysics [538]. Mid [631]. Mid-Century [631]. Midonick [716]. Mildred [261]. Milne [504]. Milton [290, 606, 628, 628]. Mind [269, 717, 154, 318, 408, 491, 693, 590, 266]. Mind-Body [154]. Minds [505]. Mirror [10]. Mises [721]. Missenard [653]. Modal [512]. Mode [577]. Model [339, 535, 493]. Models [104, 83]. Modern [232, 259, 10, 671, 587, 563, 649, 319, 86, 232, 77]. Moderne [602]. Monsarrat [511]. Montagu [177]. Montgomery [648, 131, 372]. Moon [91, 506]. Moore [248]. Moral [557, 510, 298, 361, 234, 42, 256]. Moralist [7]. Morality [85]. Morals [269]. Morgan [392]. Morris [23, 331, 90, 540]. Mortem [242]. Movement [107]. Movements [715]. Mr. [228, 245]. Mumford [268]. Munitz [606]. Murray [610]. Music [634]. My [405, 94]. Mysterious [60]. Mystic [39]. Mysticism [56].

N [197, 196, 55, 176]. Naess [635]. Nagel [625]. Naming [601]. Nathaniel [592]. Natur [24]. Natural [181, 104, 627, 22]. Naturalism [618, 116, 25]. Nature [110, 400, 331, 213, 715, 256, 639, 63, 373, 714, 130]. Necessary [502, 656]. Necessity [512]. Neglect [584]. Nelson [612, 456]. Neo [2, 252]. Neo-Kantianism [2]. Neo-Positivism [252]. Nervous [191]. Neugebauer [674]. Neumann [413]. Neurological [243]. Neurophysiology [123]. News [28, 44, 66, 96, 108, 127, 142, 157, 184, 231, 250, 285, 305, 343, 388, 433, 444, 467, 484, 495, 514, 544, 567, 579, 615, 117]. Newton [394, 688]. Nicolson [91]. Niels [672]. Ninth [369]. Noise [164]. Noland [607]. Nominal [248, 163, 249]. Non [548, 87, 49, 50]. Non-Purposeful [87, 49, 50]. Non-Quantal [548]. Nonparametric [647, 647]. Norbert [612]. Norden [158]. Norman [365, 463]. Normative [84]. Norms [482]. North [327]. Northrop [327]. Note [238, 101, 507, 341, 18, 304]. Notes [551]. Notices [28, 44, 66, 96, 108, 127, 142, 157, 184, 231, 250, 285, 305, 343, 388, 433, 444, 467, 484, 495, 514, 544, 567, 579, 615, 117]. Novak [173]. Nowell [719]. Nowell-Smith [719]. Nuclear [323]. Number [635].

O [197, 674, 407, 712]. Objectivism [302]. Objectivity [575, 247]. Obligation [719]. Occult [490]. Oehser [92]. O’Hara [160]. Oliver [304]. Oneirics [374]. Ontological [34, 476]. Ontology [338, 171, 628]. Operationism [509]. Operations [424, 192, 138, 490, 609]. Operator [476]. Opler [655]. Optimism [278]. Order [8, 304]. Organic [411]. Organism [104]. Origin [332, 252, 494]. Origins [259, 646, 413, 564]. Orman [128]. Osterburg [160]. Other [62, 505, 657, 625]. Otto [42]. Our [213]. Outlines [386].

P [110, 328, 345, 605, 57, 607, 375, 410, 430, 561, 213, 348, 716, 176, 394, 39, 719]. Pap [656, 143]. Paper [477, 227, 552, 537, 54]. Papers [459, 62]. Paradox [668, 314, 359, 669, 404, 622]. Paradoxes [392]. Parsons [9]. Parten [261]. Particle [453]. Passmore [407]. Patrick [233]. Patterns [40, 449]. Paul [717, 510, 92, 561, 407]. Paula [291]. Pavlov [657]. Peace [174]. Pears [561]. Pegasus [525]. Pei [409]. Peirce [149, 686]. Perceiving [630]. Perception [218, 266]. Perceptual [577]. Perennial [59]. Perrigo [183]. Personal [692]. Personality [492, 526]. Perspective [443, 607]. Peter [219, 254]. Peterfreund [345]. Petrovich [657]. Pettegrove [430]. Pettijohn [455]. Phenomena [664]. Phenomenalistic [3, 17]. Phenomenological [422, 577]. Philip [328, 57, 411, 483]. Philipp [77, 712, 643]. Phillips [54]. Philosopher [146]. Philosophers [665]. Philosophic [600, 574, 451, 201, 593]. Philosophic-Cosmological [574]. Philosophical [326, 161, 243, 196, 713, 223, 592, 606, 147, 254, 573, 722, 336, 503, 630, 344, 178]. Philosophie [61, 172]. Philosophy [328, 102, 527, 543, 568, 614, 605, 22, 553, 555, 301, 684, 206, 30, 317, 2, 192, 458, 587, 244, 55, 631, 634, 643, 563, 155, 246, 419, 635, 488, 423, 255, 303, 139, 313, 297, 625, 183, 620, 489, 584, 197, 198, 59, 203, 393, 561, 234, 77, 692, 176, 394, 143, 43, 463, 712, 63, 37, 329, 174, 324, 712]. Phister [648]. Physical [110, 124, 160, 213, 348, 152, 689, 447, 675, 98, 153, 712, 86, 712]. Physicalism [664]. Physicist [714]. Physics [699, 587, 672, 171, 573, 439, 1, 86, 323, 148, 457, 628, 538, 131, 634, 372]. Physik [602]. Pierre [348]. Piers [290]. Pitirim [172]. Planck [62]. Planetary [626]. Platonic [430]. Plausible [461]. Plochmann [284]. Poetry [489]. Poincaré [406]. Point [138]. Polanyi [692, 200]. Policy [365, 365]. Political [64, 76, 139]. Politics [325]. Polls [261]. Polya [461]. Pomerans [714]. Population [420]. Positivism [115, 216, 389, 252]. Possibilities [530]. Possible [535, 32]. Post [692, 242]. Post-Critical [692]. Post-Mortem [242]. Postulate [69]. Postulates [124, 488, 208]. Postwar [503]. Power [356]. Powers [58]. Pragmatism [534, 57]. Pravas [458]. Prediction [14, 403]. Predictive [524]. Prehensions [550]. Prehistory [564]. Prelude [319]. Presence [507]. Presents [500]. President [684]. Presupposition [222]. Presuppositions [399]. Pribram [40]. Priestley [629]. Primitives [239]. Principle [165, 518, 645, 33, 99, 27, 644, 667]. Principles [131, 372, 678, 451, 671, 600]. prior [283, 284]. Priori} [209]. Probability [363, 141, 522, 74, 459, 721]. Problem [144, 34, 123, 680, 139, 98, 153]. Problems [70, 225, 243, 541, 122, 573, 313, 365, 593, 414, 235, 370]. Processes [114, 611, 611]. Production [492, 296, 526]. Professor [300, 2, 141, 525, 707, 227, 170, 284, 404, 98]. Progrès [41]. Progress [330, 65]. Promise [649]. Proper [109]. Proportionally [499]. Propositions [566, 519, 84]. Protistans [702]. Psychiatry [280, 655, 620, 655]. Psycho [107]. Psycho-Linguistic [107]. Psychoanalysis [329]. Psychological [258, 82]. Psychology [589, 380, 422, 657, 448, 169, 123, 517, 426]. Psychosomatics [374]. Publication [449]. Pure [21, 441, 9]. Purpose [698]. Purposeful [87, 49, 50]. Purposefulness [48].

Q [368]. Q-Technique [368]. Qua [187]. Quality [283, 284]. Quantal [548]. Quantification [164]. Quantified [499]. Quantity [283, 284]. Quantum [114, 2, 113, 207, 548, 549, 604, 558, 1, 690, 257]. Question [140]. Quine [525, 128].

R [181, 291, 37, 363, 199, 182, 413, 94, 627, 561, 344, 176, 288]. Radcliffe [627]. Radcliffe-Brown [627]. Ragnar [482]. Ralli [410]. Ralph [59]. Raman [148]. Ramming [38]. Ramony [118]. Raoul [41]. Raphael [331]. Rapports [172]. Rather [720]. Rational [271, 537]. Rationnelle [41]. Ratner [391, 287]. Ray [632]. Raymond [681]. Readings [64, 365, 328]. Real [537]. Realism [21, 2, 18, 19, 447]. Realistic [17, 3]. Reality [194, 542, 113, 152, 98, 153, 712, 654]. Reappraisal [354, 532]. Reason [671, 331, 90]. Reasoning [336, 461]. Received [95, 111, 236, 272, 292, 306, 333, 349, 377, 376, 395, 415, 432, 466, 513, 565, 658]. Reconstruction [21]. Reconstructionism [16]. Redfield [481]. Reduction [572]. Reference [33]. Reflections [4, 643, 123, 311]. Reflexions [504]. Reformer [366]. Reformulation [188]. Reginald [556]. Reinhard [428]. Rejoinder [645, 50]. Relation [206, 597, 10, 123]. Relativism [88]. Relativity [194, 101, 298, 552, 404, 418, 314, 100, 359, 425, 547, 402, 487, 447, 504]. Religion [488, 106, 679]. Remarks [506, 603, 313]. Renaissance [430].

Reply [706, 708, 360, 405, 704, 228, 455, 362]. Report [423]. Representation [621]. Representative [722]. Representativeness [101, 100]. Research [291, 353, 424, 192, 240, 106, 441, 247, 443, 620, 258, 347, 364]. Resolution [668]. Respond [53]. Responsibility [666]. Restless [264]. Retardation [623]. Révész [564]. Review [62, 63, 64, 78, 110, 144, 175, 181, 197, 198, 216, 218, 232, 233, 258, 259, 261, 262, 263, 264, 265, 266, 267, 268, 270, 271, 289, 290, 291, 321, 325, 326, 327, 328, 132, 37, 345, 346, 347, 367, 368, 369, 161, 366, 695, 718, 129, 508, 511, 542, 588, 589, 363, 605, 22, 41, 180, 538, 553, 555, 509, 560, 651, 693, 65, 145, 199, 158, 182, 195, 426, 159, 226, 23, 196, 260, 670, 201, 118, 633, 557, 414, 713, 24, 40, 646, 717, 332, 412, 413, 91, 10, 481, 482, 650, 57, 317, 657, 25, 149, 322, 459, 541, 607, 612, 648]. Review [681, 59, 179, 203, 219, 464, 465, 510, 609, 610, 611, 58, 656, 94, 160, 177, 235, 370, 374, 677, 318, 371, 373, 375, 408, 410, 428, 429, 430, 431, 480, 591, 592, 61, 92, 131, 694, 393, 411, 608, 627, 652, 389, 458, 561, 587, 676, 60, 427, 251, 55, 606, 631, 673, 714, 329, 331, 539, 556, 632, 634, 711, 655, 147, 213, 257, 323, 330, 348, 372, 392, 563, 672, 130, 148, 234, 77, 200, 252, 254, 217, 590, 593, 653, 654, 26, 512, 715, 629, 390, 716, 710, 109, 174, 202, 214]. Review [146, 364, 635, 562, 692, 720, 128, 344, 406, 253, 286, 564, 678, 172, 38, 674, 540, 649, 558, 176, 215, 287, 324, 394, 461, 462, 624, 27, 457, 42, 320, 39, 75, 143, 255, 554, 559, 680, 721, 319, 93, 43, 173, 460, 56, 89, 76, 409, 463, 675, 679, 722, 90, 79, 407, 288, 625, 183, 256, 626, 9, 630, 178]. Reviews [269, 391, 647, 671, 365, 483, 479, 628, 712, 719]. Revised [121]. Revolution [561, 626]. Rhees [344]. Rhythmic [191]. Rice [483]. Rich [442]. Richard [718, 55, 721, 679]. Rickert [38]. Right [671, 688]. Robert [132, 588, 481, 652, 394, 173]. Roderick [630]. Roger [591, 244, 251]. Role [105]. Roles [482]. Rolf [374]. Rollo [526]. Romanell [233]. Rommetveit [482]. Roots [607]. Rose [367]. Ross [593]. Routine [212]. Rudolf [158, 512, 390, 720]. Rules [693, 499, 342]. Run [454, 401]. Runes [269, 94, 217]. Russell [266, 269, 588, 317, 365, 364, 470, 36]. Russian [197]. Ruth [429]. Ryle [300, 407].

S [327, 346, 149, 322, 647, 671, 214, 176, 324, 394, 675, 679, 626]. Sa [172]. Salmon [454]. Sambursky [675]. Saminsky [634]. Samples [261]. Samuel [654]. Santiago [118]. Sarton [267, 26]. Savage [460]. Saw [429]. Scarcity [619]. Scepticism [51]. Schaffner [611]. Scheme [532]. Schemes [281]. Schlick [63]. Schools [309, 337]. Science [232, 102, 527, 543, 568, 614, 129, 339, 166, 221, 555, 4, 279, 224, 532, 687, 475, 703, 104, 684, 312, 464, 151, 32, 582, 72, 694, 282, 627, 643, 399, 137, 452, 46, 77, 155, 246, 419, 500, 19, 488, 423, 39, 303, 106, 103, 441, 85, 449, 313, 576, 297, 712, 620, 489, 628, 259, 262, 266, 267, 328, 363, 605, 58, 92, 458, 587, 251, 26, 146, 554, 625, 132, 203, 480, 711, 330, 629, 42, 679]. Sciences [663, 642, 301, 471, 385, 302, 160, 584, 33, 635, 674, 383, 382, 647, 346, 367, 183]. Scientific [289, 346, 347, 353, 6, 317, 607, 280, 608, 606, 474, 331, 523, 601, 296, 83, 136, 451, 618, 247, 62, 363, 556, 559]. Scientist [366, 187, 7, 158]. Scientists [717]. Scope [59]. Search [650, 251, 653, 654, 75]. Sebastian [76]. Second [611, 372]. Seely [324]. Selected [542, 459]. Selections [722, 78]. Sellars [37]. Semantic [18, 164, 19]. Semantical [718]. Semantics [512, 678, 656, 255]. Semiotic [21, 53]. Sense [502]. Sense-Datum [502]. Sequential [621]. Series [330]. ses [172]. Sesonske [719]. Set [69, 342]. Seventeenth [679]. Seventeenth-Century [679]. Seventh [330]. Shame [650, 290]. Sheldon [345, 554]. Sheppard [541]. Ship [93]. Short [454, 401]. Sidney [391, 647, 287]. Siegel [647]. Significance [597]. Signs [522]. Simmel [428]. Simple [229]. Simplicity [125, 518, 645, 30, 644, 456]. Singer [290, 276, 75]. Sinnott [590]. Sir [39]. Skepticism [13, 52]. Slater [257]. Smith [719, 392]. Social [265, 367, 366, 663, 384, 301, 279, 426, 104, 717, 30, 312, 302, 584, 365, 608, 652, 33, 655, 46, 316, 296, 364, 240, 635, 383, 303, 441, 382, 482, 652]. Sociale [624]. Societies [629]. Society [166, 104, 541, 336, 365, 627, 90]. Sociological [206, 609, 31, 9, 375]. Sociologie [172]. Sociology [122, 676, 89]. Solving [144]. Some [225, 221, 437, 243, 104, 713, 385, 665, 399, 213, 46, 311, 603, 596, 313]. Somerville [174]. Son [94]. Sons [92]. Sophists [431]. Sorokin [172]. Source [322]. Soviet [574]. Space [457, 606]. Special [33, 404, 314, 359, 504]. Species [332]. Spencer [506]. Spinoza [269]. Spirit [176, 629]. Stabler [288]. Stallknecht [176, 394]. Stark [676]. State [113]. Statement [684]. Statements [398]. States [201]. Statistical [357]. Statistics [459, 647, 721, 647, 460]. Stauffer [132]. Stellung [24]. Stephan [508]. Stephen [346, 555, 159, 677, 678]. Stephenson [368]. Sterling [393]. Stewart [251]. Story [93]. Strawson [561]. Stroll [465]. Structure [557, 533, 348, 172, 517, 652]. Stuart [109, 407]. Studies [243, 240]. Study [368, 718, 363, 481, 25, 373, 694, 608, 673, 512, 109, 558, 116, 183, 630]. Styles [240]. Subjective [422]. Subjectivism [302]. Subjectivity [247]. Sullivan [280]. Superfluity [703]. Superiority [586]. Supernaturalism [116]. Superrational [271]. Support [168]. Supported [690]. Surprise [8]. Surveys [261]. Survival [448]. Svendsen [628]. Swabey [462]. Switzerland [423]. Syllogism [342, 170, 121]. Syllogisms [499]. Symbolic [23, 479, 145]. Symbolische [390]. Symmetry [536, 558]. Synonymy [358]. Synthesis [412]. Synthetic [665, 209]. System [663, 113, 441, 191, 84, 612]. Systematic [356]. Systematics [70]. Systèmes [172]. Systems [281, 355, 689, 596, 265, 161].

T [64, 326, 459, 319]. Talcott [9]. Technician [7]. Technics [268]. Technique [368]. Technology [182, 336]. Teggart [65]. Term [523]. Thales [286]. Their [159]. Themes [563]. Theodore [345, 710]. Theodosius [332]. Theoretical [549, 517]. Theories [221, 713, 6, 281, 457]. Theory [321, 325, 381, 363, 686, 180, 384, 437, 222, 338, 386, 223, 188, 105, 35, 550, 276, 242, 465, 190, 314, 341, 453, 699, 652, 257, 548, 359, 155, 246, 304, 229, 451, 558, 245, 578, 404, 504, 9, 690, 719, 110, 718, 348, 367]. There [209]. Thermodynamics [669]. Thielens [717]. Thinking [508, 214, 511]. Third [370]. Thomas [149, 626]. Thompson [210]. Thomson [321]. Thought [346, 542, 589, 201, 149, 607, 574, 86, 64, 40, 646, 556, 288, 626]. Thoughts [715]. Three [713]. Time [124, 230, 717, 506, 606, 295, 472, 138, 123, 229]. Times [504]. Tobias [406]. Tomorrow [610]. Topitsch [537]. Törnebohm [405]. Toulmin [555, 677]. Tradition [345, 422]. Transactions [369, 611, 370]. Transition [319]. Translation [225]. Translations [254]. Transplanar [211]. Treatment [522]. Trends [503]. Trevett [556]. True [715]. Truth [685, 297, 721, 718]. Turth [656]. Two [278, 240, 689, 504, 219]. Types [531].

Ullman [678]. Ullmann [159]. Understanding [676]. Unextended [167]. Unfinished [447]. Unfreedom [421]. Unitary [99]. unité [41]. United [201]. Unity [385]. Universal [506, 251, 472]. Universe [264, 60]. Unreason [90]. Untersteiner [431]. Use [400, 464, 159]. Uses [677].

V [37, 148]. Validity [499]. Valuation [25, 474]. Value [437, 71, 386, 223, 151, 582, 282, 483, 361, 83, 187, 719, 539, 632, 540]. Values [223, 655, 365]. Varieties [540]. Veatch [260]. Vectorial [425]. Velocities [425]. Vergilius [161]. Verifiability [125, 551, 644, 116]. Vernacular [598]. Vers [41]. versus [3, 17, 700, 152]. Very [442]. Victor [252]. Vienna [252]. View [494, 649]. Viewpoints [481]. Views [713]. Violence [139]. Virchow [720]. Viscount [654]. Vision [287]. Visual [218]. Volume [12, 67, 81, 119, 134, 185, 205, 274, 350, 434, 496, 569, 636, 723, 635]. Voyages [91].

W [110, 266, 37, 511, 459, 160, 671, 131, 561, 483, 147, 213, 323, 372, 593, 629, 649, 407]. W. [304]. Wagner [717]. Wald [459]. Walter [628]. War [424]. Warburg [494]. Ware [590]. Warnock [561]. Was [688]. Watkins [325]. Watson [131, 372, 373]. Way [75]. Web [428]. Weltbild [24]. Wentworth [463]. Werkmeister [226, 147]. Werner [676, 714]. Wertphilosophie [38]. Weschler [291]. West [317, 31]. Western [217, 176, 626]. Westfall [679]. Whether [702]. White [129, 694]. Whitehead [327, 507, 550, 15, 402, 489, 592]. Whiteheadian [190]. Whittaker [79]. Whole [481]. Whorf [542]. Wiener [607, 328, 57, 348]. Wiezsäcker [130]. Wilcox [199]. Will [530]. Willard [128]. William [368, 131, 372, 407]. Williams [591, 320]. Wilson [347]. Wisdom [329]. Wissenschaft [24]. without [239, 625]. Witt [463]. Wittgenstein [344]. Wolff [428]. Wollheim [561]. Wolstein [25]. Words [159]. Wordsworth [489]. Work [681]. World [218, 651, 675, 86]. Wright [722, 169, 680]. Writings [542, 254, 722, 326].

Yalden [321]. Yalden-Thomson [321].

Ziff [407]. Zipf [27]. Zuce [144]. Zurich [423].

References

Werkmeister:1950:EBQ

[1] W. H. Werkmeister. An epistemological basis for quantum physics. Philosophy of Science, 17(1):1–25, January 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185673.

Grunbaum:1950:RNK

[2] Adolf Grünbaum. Realism and neo-Kantianism in Professor Margenau’s philosophy of quantum mechanics. Philosophy of Science, 17(1):26–34, January 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185674.

Feigl:1950:EHR

[3] Herbert Feigl. Existential hypotheses. Realistic versus phenomenalistic interpretations. Philosophy of Science, 17(1):35–62, January 1950. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185675.

Bridgman:1950:IRH

[4] P. W. Bridgman. Impertinent reflections on history of science. Philosophy of Science, 17(1):63–73, January 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185676.

Beck:1950:CIE

[5] Lewis White Beck. Constructions and inferred entities. Philosophy of Science, 17(1):74–86, January 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185677.

Freund:1950:CST

[6] John E. Freund. On the confirmation of scientific theories. Philosophy of Science, 17(1):87–94, January 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185678.

Simpson:1950:STM

[7] George Simpson. The scientist — technician or moralist? Philosophy of Science, 17(1):95–108, January 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185679.

Gardner:1950:OS

[8] Martin Gardner. Order and surprise. Philosophy of Science, 17(1):109– 117, January 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185680.

Williams:1950:BRB

[9] Richard Hays Williams. Book review: Essays in Sociological Theory: Pure and Applied by Talcott Parsons. Philosophy of Science, 17(1):118–119, January 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185681.

Foote:1950:BRB

[10] Nelson N. Foote. Book review: Mirror for Man: The Relation of Anthropology to Modern Life by Clyde Kluckhohn. Philosophy of Science, 17 (1):120–121, January 1950. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185682.

Anonymous:1950:FMa

[11] Anonymous. Front matter. Philosophy of Science, 17(1):??, January 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185672.

Anonymous:1950:VIa

[12] Anonymous. Volume information. Philosophy of Science, 17(1):??, January 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/185671.

Berenda:1950:FFS

[13] Carlton W. Berenda. A five-fold skepticism in logical empiricism. Philosophy of Science, 17(2):123–132, April 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184911.

Miller:1950:BDP

[14] David L. Miller. The behavioral dimension of prediction and meaning. Philosophy of Science, 17(2):133–141, April 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184912.

Lawrence:1950:WME

[15] Nathaniel Lawrence. Whitehead’s method of extensive abstraction. Philosophy of Science, 17(2):142–163, April 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184913.

Churchman:1950:LR

[16] C. West Churchman. Logical reconstructionism. Philosophy of Science, 17(2):164–166, April 1950. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184914.

Frank:1950:CRV

[17] Philipp Frank. Comments on realistic versus phenomenalistic interpretations. Philosophy of Science, 17(2):166–169, April 1950. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184915.

Hempel:1950:NSR

[18] Carl G. Hempel. A note on semantic realism. Philosophy of Science, 17(2):169–173, April 1950. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184916.

Nagel:1950:SSR

[19] Ernest Nagel. Science and semantic realism. Philosophy of Science, 17(2): 174–181, April 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184917.

Ramsperger:1950:FEH

[20] A. G. Ramsperger. On Feigl’s “existential hypotheses”. Philosophy of Science, 17(2):182–185, April 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184918.

Feigl:1950:LRR

[21] Herbert Feigl. Logical reconstruction, realism and pure semiotic. Philosophy of Science, 17(2):186–195, April 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184919.

Bergmann:1950:BRB

[22] Gustav Bergmann. Book review: Natural Philosophy of Cause and Chance by Max Born. Philosophy of Science, 17(2):196–199, April 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184920.

Copi:1950:BRB

[23] Irving M. Copi. Book review: Fundamentals of Symbolic Logic by Alice Ambrose; Morris Lazerowitz. Philosophy of Science, 17(2):199–200, April 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184921.

Dornfeld:1950:BRB

[24] Ernst J. Dornfeld. Book review: Das Biologische Weltbild. Band I. Die Stellung des Lebens in Natur und Wissenschaft by Ludwig von Bertalanffy. Philosophy of Science, 17(2):200–202, April 1950. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184922.

Geiger:1950:BRB

[25] George R. Geiger. Book review: Experience and Valuation: A Study in John Dewey’s Naturalism by Benjamin Wolstein. Philosophy of Science, 17(2):202–203, April 1950. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184923.

Lundberg:1950:BRB

[26] George A. Lundberg. Book review: The Life of Science by George Sarton. Philosophy of Science, 17(2):203–204, April 1950. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184924.

Riemer:1950:BRB

[27] Svend Riemer. Book review: Human Behavior and the Principle of Least Effort. An Introduction to Human Ecology by George K. Zipf. Philosophy of Science, 17(2):204–205, April 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184925.

Anonymous:1950:NNa

[28] Anonymous. News and notices. Philosophy of Science, 17(2):206, April 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184926.

Anonymous:1950:FMb

[29] Anonymous. Front matter. Philosophy of Science, 17(2):??, April 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184910.

Fries:1950:LSC

[30] Horace S. Fries. Logical simplicity: A challenge to philosophy and to social inquiry. Philosophy of Science, 17(3):207–228, July 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185218.

Hartung:1950:SEM

[31] Frank E. Hartung. A sociological evaluation of the meeting of East and West. Philosophy of Science, 17(3):229–237, July 1950. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185219.

Hartman:1950:SEP

[32] Robert S. Hartman. Is a science of ethics possible. Philosophy of Science, 17(3):238–246, July 1950. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185220.

Henry:1950:PLS

[33] Jules Henry. The principle of limits: With special reference to the social sciences. Philosophy of Science, 17(3):247–253, July 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185221.

Feibleman:1950:CMO

[34] James K. Feibleman. Class-membership and the ontological problem. Philosophy of Science, 17(3):254–259, July 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185222.

Fitch:1950:TFH

[35] Frederic B. Fitch and Gladys Barry. Towards a formalization of Hull’s behavior theory. Philosophy of Science, 17(3):260–265, July 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185223.

Hay:1950:BRJ

[36] W. H. Hay. Bertrand Russell on the justification of induction. Philosophy of Science, 17(3):266–277, July 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185224.

Ackoff:1950:BRB

[37] Russell L. Ackoff. Book review: Philosophy for the Future by R. W. Sellars; V. J. McGill; M. Farber. Philosophy of Science, 17(3):278–279, July 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185225.

Orden:1950:BRB

[38] Marianne Orden. Book review: Karl Jasper und Heinrich Rickert: Existenzialismus und Wertphilosophie by Gustav Ramming. Philosophy of Science, 17(3):279, July 1950. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185226.

Ruddick:1950:BRB

[39] C. T. Ruddick. Book review: Sir Arthur Eddington: Man of Science and Mystic by L. P. Jacks. Philosophy of Science, 17(3):279–280, July 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185227.

Dunham:1950:BRB

[40] Barrows Dunham. Book review: Conflicting Patterns of Thought by Karl Pribram. Philosophy of Science, 17(3):280–282, July 1950. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185228.

Biser:1950:BRB

[41] Erwin Biser. Book review: De Descartes à Ampère Ou Progrès Vers L’unité Rationnelle by Raoul Ferrier. Philosophy of Science, 17(3):282– 283, July 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185229.

Rothman:1950:BRB

[42] Robert Rothman. Book review: Science and the Moral Life by Max C. Otto. Philosophy of Science, 17(3):283–284, July 1950. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185230.

Sellars:1950:BRB

[43] Roy Wood Sellars. Book review: Language and Philosophy by Max Black. Philosophy of Science, 17(3):284–285, July 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185231.

Anonymous:1950:NNb

[44] Anonymous. News and notices. Philosophy of Science, 17(3):286, July 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185232.

Anonymous:1950:FMc

[45] Anonymous. Front matter. Philosophy of Science, 17(3):??, July 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185217.

Levy:1950:SBM

[46] Marion J. Levy, Jr. Some basic methodological difficulties in social science. Philosophy of Science, 17(4):287–301, October 1950. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185928.

Leblanc:1950:D

[47] Hugues Leblanc. On definitions. Philosophy of Science, 17(4):302–309, October 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185929.

Taylor:1950:CMC

[48] Richard Taylor. Comments on a mechanistic conception of purposefulness. Philosophy of Science, 17(4):310–317, October 1950. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185930.

Rosenblueth:1950:PNP

[49] Arturo Rosenblueth and Norbert Wiener. Purposeful and non-purposeful behavior. Philosophy of Science, 17(4):318–326, October 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185931.

Taylor:1950:PNP

[50] Richard Taylor. Purposeful and non-purposeful behavior: A rejoinder. Philosophy of Science, 17(4):327–332, October 1950. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185932.

Wang:1950:SAI

[51] Hao Wang. On scepticism about induction. Philosophy of Science, 17 (4):333–335, October 1950. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185933.

Will:1950:SF

[52] Frederick L. Will. Skepticism and the future. Philosophy of Science, 17 (4):336–346, October 1950. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185934.

Phillips:1950:CDR

[53] Jean Phillips. The concept “disposition to respond” in a behavioral semiotic. Philosophy of Science, 17(4):347–353, October 1950. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185935. See comment [54].

Morris:1950:CPJ

[54] Charles Morris. Comment on the paper by Jean A. Phillips. Philosophy of Science, 17(4):354–355, October 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 185936. See [53].

Isenberg:1950:BRB

[55] Arnold Isenberg. Book review: A Philosophy of Life by Richard N. Bender. Philosophy of Science, 17(4):356, October 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185937.

Sheldon:1950:BRB

[56] W. H. Sheldon. Book review: Introduction to Comparative Mysticism by Jacques de Marquette. Philosophy of Science, 17(4):356, October 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185938.

Fries:1950:BRB

[57] Horace Fries. Book review: Evolution and the Founders of Pragmatism by Philip P. Wiener. Philosophy of Science, 17(4):357, October 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185939.

Goran:1950:BRB

[58] Morris Goran. Book review: The Powers and Limits of Science by E. F. Caldin. Philosophy of Science, 17(4):358, October 1950. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185940.

Golightly:1950:BRB

[59] Cornelius L. Golightly. Book review: The Perennial Scope of Philosophy by Karl Jaspers; Ralph Manheim. Philosophy of Science, 17(4):358–359, October 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185941.

Hirsch:1950:BRB

[60] Helmut Hirsch. Book review: Man and This Mysterious Universe by Brynjolf Bjorset. Philosophy of Science, 17(4):359–360, October 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185942.

Hartung:1950:BRB

[61] Frank E. Hartung. Book review: Europäische Philosophie der Gegenwart by I. M. Bochenski. Philosophy of Science, 17(4):360–361, October 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185943.

A:1950:BRBa

[62] R. L. A. Book review: Scientific Autobiography and Other Papers by Max Planck. Philosophy of Science, 17(4):361, October 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185944.

A:1950:BRBb

[63] R. L. A. Book review: Philosophy of Nature by M. Schlick. Philosophy of Science, 17(4):361, October 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185945.

A:1950:BRBc

[64] R. L. A. Book review: Free Government in the Making: Readings in American Political Thought by A. T. Mason. Philosophy of Science, 17 (4):361–362, October 1950. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185946.

C:1950:BRB

[65] C. W. C. Book review: The Idea of Progress by F. J. Teggart; G. H. Hildebrand. Philosophy of Science, 17(4):362, October 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185947.

Anonymous:1950:NNc

[66] Anonymous. News and notices. Philosophy of Science, 17(4):363, October 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185948.

Anonymous:1950:VIb

[67] Anonymous. Volume information. Philosophy of Science, 17(4):365–366, October 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185926.

Anonymous:1950:FMd

[68] Anonymous. Front matter. Philosophy of Science, 17(4):??, October 1950. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185927.

Cowan:1951:PSE

[69] Thomas A. Cowan. A postulate set for experimental jurisprudence. Philosophy of Science, 18(1):1–15, January 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185875.

Alderson:1951:SPA

[70] Wroe Alderson. A systematics for problems of action. Philosophy of Science, 18(1):16–25, January 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185876.

Cerf:1951:VD

[71] Walter Cerf. Value decisions. Philosophy of Science, 18(1):26–34, January 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185877.

Hartung:1951:SI

[72] Frank E. Hartung. Science as an institution. Philosophy of Science, 18 (1):35–54, January 1951. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185878.

Weakland:1951:MCA

[73] John Hast Weakland. Method in cultural anthropology. Philosophy of Science, 18(1):55–69, January 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185879.

Martin:1951:ECC

[74] Norman M. Martin. The explicandum of the classical concept of probability. Philosophy of Science, 18(1):70–84, January 1951. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185880.

Ruddick:1951:BRB

[75] C. T. Ruddick. Book review: In Search of a Way of Life by Edgar Arthur Singer. Philosophy of Science, 18(1):85–86, January 1951. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185881.

Simpson:1951:BRB

[76] George Simpson. Book review: The Political Community by Sebastian De Grazia. Philosophy of Science, 18(1):86–87, January 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185882.

Lindsay:1951:BRB

[77] R. B. Lindsay. Book review: Modern Science and Its Philosophy by Philipp Frank. Philosophy of Science, 18(1):87–88, January 1951. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185883.

A:1951:BRBa

[78] R. L. A. Book review: Selections from Bergson by H. A. Larrabee. Philosophy of Science, 18(1):88, January 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185884.

Struik:1951:BRB

[79] D. J. Struik. Book review: From Euclid to Eddington by E. Whittaker. Philosophy of Science, 18(1):88–91, January 1951. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185885.

Anonymous:1951:FMa

[80] Anonymous. Front matter. Philosophy of Science, 18(1):??, January 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185874.

Anonymous:1951:VIa

[81] Anonymous. Volume information. Philosophy of Science, 18(1):??, January 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/185873.

Bergmann:1951:LPC

[82] Gustav Bergmann. The logic of psychological concepts. Philosophy of Science, 18(2):93–110, April 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185656.

Meyer:1951:HVS

[83] Herman Meyer. On the heuristic value of scientific models. Philosophy of Science, 18(2):111–123, April 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185657.

Werkmeister:1951:NPI

[84] W. H. Werkmeister. Normative propositions and the ideal of an integrated and closed system. Philosophy of Science, 18(2):124–131, April 1951. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185658.

Simpson:1951:SM

[85] George Simpson. Science as morality. Philosophy of Science, 18(2):132– 143, April 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185659.

Yaker:1951:MTM

[86] Henri Marc Yaker. Medieval thought, modern physics, and the physical world. Philosophy of Science, 18(2):144–153, April 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185660.

Moulyn:1951:PNP

[87] Adrian C. Moulyn. Purposeful and non-purposeful behavior. Philosophy of Science, 18(2):154, April 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185661.

Will:1951:REI

[88] Frederick L. Will. Relativism and experimental inference. Philosophy of Science, 18(2):155–169, April 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185662.

Sheppard:1951:BRB

[89] Harold Sheppard. Book review: An Introduction to the History of Sociology by Harry Elmer Barnes. Philosophy of Science, 18(2):170–171, April 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185663.

Strong:1951:BRB

[90] Samuel M. Strong. Book review: Reason and Unreason in Society by Morris Ginsberg. Philosophy of Science, 18(2):171–172, April 1951. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185664.

Finch:1951:BRB

[91] Herbert H. Finch. Book review: Voyages to the Moon by Marjorie Hope Nicolson. Philosophy of Science, 18(2):172–173, April 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185665.

Hartung:1951:BRB

[92] Rosemary Hartung. Book review: Sons of Science by Paul H. Oehser. Philosophy of Science, 18(2):173, April 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185666.

Schmidt:1951:BRB

[93] Lester F. Schmidt. Book review: The Story of the Ship by Charles E. Gibson. Philosophy of Science, 18(2):173–174, April 1951. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185667.

H:1951:BRB

[94] F. H. Book review: Letters to My Son by Dagobert R. Runes. Philosophy of Science, 18(2):174, April 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185668.

Anonymous:1951:BRa

[95] Anonymous. Books received. Philosophy of Science, 18(2):175, April 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185669.

Anonymous:1951:NNa

[96] Anonymous. News and notices. Philosophy of Science, 18(2):176–181, April 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/185670.

Anonymous:1951:FMb

[97] Anonymous. Front matter. Philosophy of Science, 18(2):??, April 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185655.

Werkmeister:1951:PMP

[98] W. H. Werkmeister. Professor Margenau and the problem of physical reality. Philosophy of Science, 18(3):183–192, July 1951. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185700.

Lillie:1951:BUP

[99] Ralph S. Lillie. Biology and unitary principle. Philosophy of Science, 18(3):193–207, July 1951. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185701.

Hammond:1951:RR

[100] Kenneth R. Hammond. Relativity and representativeness. Philosophy of Science, 18(3):208–211, July 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 185702. See note [101].

Brunswik:1951:NHA

[101] Egon Brunswik. Note on Hammond’s analogy between “Relativity and Representativeness”. Philosophy of Science, 18(3):212–217, July 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185703. See [100].

Adams:1951:CPP

[102] John Stokes Adams, Jr. Contemporary philosophy and philosophy of science. Philosophy of Science, 18(3):218–222, July 1951. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185704.

Searles:1951:FCE

[103] Herbert L. Searles. Form and content in empirical science. Philosophy of Science, 18(3):223–229, July 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185705.

Deutsch:1951:MOS

[104] Karl W. Deutsch. Mechanism, organism, and society: Some models in natural and social science. Philosophy of Science, 18(3):230–252, July 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185706.

Fiala:1951:ERT

[105] Silvio Fiala. The experiment and its role in the theory of knowledge. Philosophy of Science, 18(3):253–258, July 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185707.

Rusk:1951:CRS

[106] George Y. Rusk. Cooperative research in the science of religion. Philosophy of Science, 18(3):259–261, July 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185708.

Vuysje:1951:PLM

[107] D. Vuysje. The psycho-linguistic movement in Holland. Philosophy of Science, 18(3):262–268, July 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185709.

Anonymous:1951:NNb

[108] Anonymous. News and notices. Philosophy of Science, 18(3):269, July 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185710.

McCloskey:1951:BRB

[109] Burr McCloskey. Book review: The Proper Study of Mankind by Stuart Chase. Philosophy of Science, 18(3):270–271, July 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185711.

A:1951:BRBb

[110] R. L. A. Book review: The Nature of Physical Theory by P. W. Bridgman. Philosophy of Science, 18(3):271, July 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185712.

Anonymous:1951:BRb

[111] Anonymous. Books received. Philosophy of Science, 18(3):272, July 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185713.

Anonymous:1951:FMc

[112] Anonymous. Front matter. Philosophy of Science, 18(3):??, July 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185699.

Kemble:1951:RMS

[113] Edwin C. Kemble. Reality, measurement, and the state of the system in quantum mechanics. Philosophy of Science, 18(4):273–299, October 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185563. See comment [227] and reply

[228].

Elsasser:1951:QMA

[114] Walter M. Elsasser. Quantum mechanics, amplifying processes, and living matter. Philosophy of Science, 18(4):300–326, October 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185564.

Cerf:1951:LPE

[115] Walter Cerf. Logical positivism and existentialism. Philosophy of Science, 18(4):327–338, October 1951. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185565.

Spiegelberg:1951:SNS

[116] Herbert Spiegelberg. Supernaturalism or naturalism: A study in meaning and verifiability. Philosophy of Science, 18(4):339–368, October 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185566.

Churchman:1951:NN

[117] C. West Churchman. News and notices. Philosophy of Science, 18(4):369, October 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185567.

Cragie:1951:BRB

[118] E. H. Cragie. Book review: Explorer of the Human Brain: The Life of Santiago Ramony Cajal by Dorothy F. Cannon. Philosophy of Science, 18(4):370, October 1951. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185568.

Anonymous:1951:VIb

[119] Anonymous. Volume information. Philosophy of Science, 18(4):371–372, October 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185561.

Anonymous:1951:FMd

[120] Anonymous. Front matter. Philosophy of Science, 18(4):??, October 1951. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185562.

Reichenbach:1952:SR

[121] Hans Reichenbach. The syllogism revised. Philosophy of Science, 19(1):1– 16, January 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185095.

Hartung:1952:PSK

[122] Frank E. Hartung. Problems of the sociology of knowledge. Philosophy of Science, 19(1):17–32, January 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185096.

Moulyn:1952:RPT

[123] Adrian C. Moulyn. Reflections on the problem of time in relation to neurophysiology and psychology. Philosophy of Science, 19(1):33–49, January 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185097.

Biser:1952:PPT

[124] Erwin Biser. Postulates for physical time. Philosophy of Science, 19(1):50– 69, January 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185098.

Berenda:1952:VSE

[125] C. W. Berenda. On verifiability, simplicity, and equivalence. Philosophy of Science, 19(1):70–76, January 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185099.

Storer:1952:LI

[126] Thomas Storer. Linguistic isomorphisms. Philosophy of Science, 19(1):77– 85, January 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185100.

Anonymous:1952:NNa

[127] Anonymous. News and notices. Philosophy of Science, 19(1):86, January 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185101.

Nakhnikian:1952:BRB

[128] George Nakhnikian. Book review: Methods of Logic by Willard Van Orman Quine. Philosophy of Science, 19(1):87, January 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185102.

Barnes:1952:BRB

[129] Harry Elmer Barnes. Book review: The Science of Culture by Leslie A. White. Philosophy of Science, 19(1):87–89, January 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185103.

Layzer:1952:BRBa

[130] David Layzer. Book review: The History of Nature by C. F. von Wiezsäcker. Philosophy of Science, 19(1):89–90, January 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185104.

Hartung:1952:BRB

[131] Frank Hartung. Book review: Physics: Principles and Applications by Henry Margenau; William W. Watson; Carol G. Montgomery. Philosophy of Science, 19(1):90–91, January 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185105.

Ackley:1952:BRB

[132] Sheldon Ackley. Book review: Science and Civilization by Robert C. Stauffer. Philosophy of Science, 19(1):91–92, January 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185106.

Anonymous:1952:FMa

[133] Anonymous. Front matter. Philosophy of Science, 19(1):??, January 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185094.

Anonymous:1952:VIa

[134] Anonymous. Volume information. Philosophy of Science, 19(1):??, January 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/185093.

Reiser:1952:EC

[135] Oliver L. Reiser. The evolution of cosmologies. Philosophy of Science, 19(2):93–107, April 1952. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185818.

Munitz:1952:SMC

[136] Milton K. Munitz. Scientific method in cosmology. Philosophy of Science, 19(2):108–130, April 1952. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185819.

King:1952:MES

[137] Lester S. King. Is medicine an exact science? Philosophy of Science, 19(2):131–140, April 1952. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185820.

Moulyn:1952:FPL

[138] Adrian C. Moulyn. The functions of point and line in time measuring operations. Philosophy of Science, 19(2):141–155, April 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185821.

Somerville:1952:KPC

[139] John Somerville. A key problem of current political philosophy: The issue of force and violence. Philosophy of Science, 19(2):156–165, April 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185822.

Fries:1952:FDC

[140] Horace S. Fries. Is the Ψ-function description “complete?” A layman’s question. Philosophy of Science, 19(2):166–169, April 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185823.

Hay:1952:PCP

[141] William H. Hay. Professor Carnap and probability. Philosophy of Science, 19(2):170–177, April 1952. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185824.

Anonymous:1952:NNb

[142] Anonymous. News and notices. Philosophy of Science, 19(2):178, April 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185825.

Ruddick:1952:BRB

[143] C. T. Ruddick. Book review: Elements of Analytic Philosophy by Arthur Pap. Philosophy of Science, 19(2):179, April 1952. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185826.

A:1952:BRBa

[144] R. L. A. Book review: Essentials in Problem Solving by Zuce Kogan. Philosophy of Science, 19(2):180, April 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185827.

C:1952:BRB

[145] C. W. C. Book review: Symbolic Logic by C. I. Lewis; C. H. Langford. Philosophy of Science, 19(2):180, April 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185828.

McIntyre:1952:BRB

[146] Henry C. McIntyre. Book review: Francis Bacon: Philosopher of Industrial Science by Benjamin Farrington. Philosophy of Science, 19(2):180, April 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/185829.

Klausner:1952:BRB

[147] Neal W. Klausner. Book review: A History of Philosophical Ideas in America by W. H. Werkmeister. Philosophy of Science, 19(2):180–181, April 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/185830.

Layzer:1952:BRBb

[148] David Layzer. Book review: The New Physics by C. V. Raman. Philosophy of Science, 19(2):181–182, April 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185831.

Geiger:1952:BRB

[149] George R. Geiger. Book review: The Thought of C. S. Peirce by Thomas A. Goudge. Philosophy of Science, 19(2):182, April 1952. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185832.

Anonymous:1952:FMb

[150] Anonymous. Front matter. Philosophy of Science, 19(2):??, April 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185817.

Gotshalk:1952:VS

[151] D. W. Gotshalk. Value science. Philosophy of Science, 19(3):183–192, July 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185728.

Margenau:1952:PVH

[152] Henry Margenau. Physical versus historical reality. Philosophy of Science, 19(3):193–213, July 1952. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185729.

Werkmeister:1952:PPR

[153] W. H. Werkmeister. The problem of physical reality. Philosophy of Science, 19(3):214–224, July 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185730.

Golightly:1952:MBC

[154] Cornelius L. Golightly. Mind-body, causation and correlation. Philosophy of Science, 19(3):225–227, July 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185731.

Madden:1952:PSG

[155] Edward H. Madden. The philosophy of science in gestalt theory. Philosophy of Science, 19(3):228–238, July 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185732.

McKinney:1952:D

[156] J. P. McKinney. Discussion. Philosophy of Science, 19(3):239–241, July 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185733.

Anonymous:1952:NNc

[157] Anonymous. News and notices. Philosophy of Science, 19(3):242, July 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185734.

Cappannari:1952:BRBa

[158] Stephen C. Cappannari. Book review: Goethe as a Scientist by Rudolf Magnus; Heinz Norden. Philosophy of Science, 19(3):243, July 1952. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185737.

Collinson:1952:BRB

[159] John Collinson. Book review: Words and Their Use by Stephen Ullmann. Philosophy of Science, 19(3):243, July 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185735.

H:1952:BRBa

[160] F. E. H. Book review: An Introduction to Criminalistics: The Application of the Physical Sciences to the Detection of Crime by Charles E. O’Hara; James W. Osterburg. Philosophy of Science, 19(3):243, July 1952. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185736.

Adams:1952:BRB

[161] John Stokes Adams, Jr. Book review: A History of Philosophical Systems by Vergilius Ferm. Philosophy of Science, 19(3):244, July 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185738.

Anonymous:1952:FMc

[162] Anonymous. Front matter. Philosophy of Science, 19(3):??, July 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185727.

Moore:1952:NDC

[163] Omar Khayyam Moore. Nominal definitions of ’culture’. Philosophy of Science, 19(4):245–256, October 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185391.

Hockett:1952:AQS

[164] Charles F. Hockett. An approach to the quantification of semantic noise. Philosophy of Science, 19(4):257–260, October 1952. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185392.

Bachem:1952:HIP

[165] A. Bachem. Heisenberg’s indeterminacy principle and life. Philosophy of Science, 19(4):261–272, October 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185393.

Becker:1952:SCS

[166] Howard Becker. Science, culture, and society. Philosophy of Science, 19 (4):273–287, October 1952. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185394.

Grunbaum:1952:CCE

[167] Adolf Grünbaum. A consistent conception of the extended linear continuum as an aggregate of unextended elements. Philosophy of Science, 19 (4):288–306, October 1952. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185395.

Kemeny:1952:DFS

[168] John G. Kemeny and Paul Oppenheim. Degree of factual support. Philosophy of Science, 19(4):307–324, October 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185396.

Madden:1952:CWL

[169] Edward H. Madden and Marian C. Madden. Chauncey Wright and the logic of psychology. Philosophy of Science, 19(4):325–332, October 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185397.

Plochmann:1952:PHF

[170] George Kimball Plochmann. Professor Henle on the four figures of syllogism. Philosophy of Science, 19(4):333–341, October 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185398.

Margenau:1952:PO

[171] Henry Margenau. Physics and ontology. Philosophy of Science, 19(4):342– 345, October 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185399.

Nisbet:1952:BRB

[172] R. A. Nisbet. Book review: Sociologie de la Connaissance, Sa Structure et ses Rapports Avec la Philosophie de la Connaissance. Étude Critique des Systèmes de Karl Mannheim et de Pitirim A. Sorokin by Jacques J. Maquet. Philosophy of Science, 19(4):346–347, October 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185400.

Sellers:1952:BRB

[173] Charles Sellers. Book review: Autiobiography of Dr. Robert Meyer by Robert Meyer; Emil Novak. Philosophy of Science, 19(4):347, October 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185401.

McCloskey:1952:BRB

[174] Burr McCloskey. Book review: Philosophy of Peace by John Somerville. Philosophy of Science, 19(4):347–348, October 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/185402.

A:1952:BRBb

[175] R. L. A. Book review: The “Why” of Man’s Experience by Hadley Cantrill. Philosophy of Science, 19(4):348, October 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185403.

Portnoy:1952:BRB

[176] Julius Portnoy. Book review: The Spirit of Western Philosophy by N. P. Stallknecht; R. S. Brumbaugh. Philosophy of Science, 19(4):348, October 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185404.

H:1952:BRBb

[177] F. E. H. Book review: On Being Human by Ashley Montagu. Philosophy of Science, 19(4):348–349, October 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185405.

Zerby:1952:BRB

[178] Lewis K. Zerby. Book review: Philosophical Analysis by Max Black. Philosophy of Science, 19(4):349, October 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185406.

Golightly:1952:BRB

[179] Cornelius L. Golightly. Book review: Cosmology by E. Finlay-Freundlich. Philosophy of Science, 19(4):349–350, October 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/185407.

Biser:1952:BRB

[180] Erwin Biser. Book review: A New Theory of Gravitation by Jakob Mandelker. Philosophy of Science, 19(4):350, October 1952. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185408.

A:1952:BRBc

[181] R. L. A. Book review: Creative Aspects of Natural Law by R. A. Fisher. Philosophy of Science, 19(4):350–351, October 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185409.

Cappannari:1952:BRBb

[182] Stephen C. Cappannari. Book review: Man the Maker: A History of Technology and Engineering by R. J. Forbes. Philosophy of Science, 19 (4):351, October 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185410.

Werkmeister:1952:BRB

[183] W. H. Werkmeister. Book review: Epitomization, A Study in Philosophy of the Sciences by George Perrigo Conger. Philosophy of Science, 19 (4):351–355, October 1952. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185411.

See comment [226].

Anonymous:1952:NNd

[184] Anonymous. News and notices. Philosophy of Science, 19(4):356, October 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185412.

Anonymous:1952:VIb

[185] Anonymous. Volume information. Philosophy of Science, 19(4):357–358, October 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185389.

Anonymous:1952:FMd

[186] Anonymous. Front matter. Philosophy of Science, 19(4):??, October 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185390.

Rudner:1953:SQS

[187] Richard Rudner. The scientist qua scientist makes value judgments. Philosophy of Science, 20(1):1–6, January 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185617.

Elsasser:1953:RBT

[188] Walter M. Elsasser. A reformulation of Bergson’s theory of memory. Philosophy of Science, 20(1):7–21, January 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185618.

Bain:1953:WC

[189] Read Bain. What is this crisis? Philosophy of Science, 20(1):22–30, January 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185619.

Greenman:1953:WTM

[190] Martin A. Greenman. A Whiteheadian theory of meaning. Philosophy of Science, 20(1):31–41, January 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185620.

Teitelbaum:1953:RAN

[191] Harry A. Teitelbaum. The rhythmic activity of the nervous system. Philosophy of Science, 20(1):42–58, January 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185621.

Hindrichs:1953:TPO

[192] Gerard Hindrichs. Toward a philosophy of operations research. Philosophy of Science, 20(1):59–66, January 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185622.

Fales:1953:CE

[193] Walter Fales. Causes and effects. Philosophy of Science, 20(1):67–74, January 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185623.

Bachem:1953:RR

[194] A. Bachem. The relativity of reality. Philosophy of Science, 20(1):75–78, January 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185624.

Cohn:1953:BRB

[195] Harvey Cohn. Book review: Johannes Kepler: Life and Letters by Carola Baumgart. Philosophy of Science, 20(1):79, January 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185625.

Copi:1953:BRBa

[196] Irving M. Copi. Book review: Logic and Language: A Collection of Philosophical Articles by A. G. N. Flew. Philosophy of Science, 20(1):79–80, January 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185626.

A:1953:BRBa

[197] R. L. A. Book review: History of Russian Philosophy by N. O. Lossky. Philosophy of Science, 20(1):80, January 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185627.

A:1953:BRBb

[198] R. L. A. Book review: A History of Ancient and Medieval Philosophy by Frederick Mayer. Philosophy of Science, 20(1):80–81, January 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185628.

C:1953:BRB

[199] C. W. C. Book review: The Anatomy of Mathematics by R. B. Kershner; L. R. Wilcox. Philosophy of Science, 20(1):81, January 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185629.

Linsky:1953:BRBa

[200] Leonard Linsky. Book review: The Logic of Liberty by Michael Polanyi. Philosophy of Science, 20(1):81, January 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185630.

Cowan:1953:BRB

[201] Thomas Cowan. Book review: Philosophic Thought in France and the United States by Marvin Farber. Philosophy of Science, 20(1):81–82, January 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/185631.

McGary:1953:BRBa

[202] Keith McGary. Book review: The Human Community by Baker Brownell. Philosophy of Science, 20(1):82, January 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185632.

Golightly:1953:BRBa

[203] Cornelius L. Golightly. Book review: Science: Its Method and Its Philosophy by G. Burniston Brown. Philosophy of Science, 20(1):83, January 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185633.

Anonymous:1953:FMa

[204] Anonymous. Front matter. Philosophy of Science, 20(1):??, January 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185616.

Anonymous:1953:VIa

[205] Anonymous. Volume information. Philosophy of Science, 20(1):??, January 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/185615.

Feuer:1953:SAR

[206] Lewis S. Feuer. Sociological aspects of the relation between language and philosophy. Philosophy of Science, 20(2):85–100, April 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185445.

Lande:1953:CKQ

[207] Alfred Landé. Continuity, a key to quantum mechanics. Philosophy of Science, 20(2):101–109, April 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185446.

Stephenson:1953:PB

[208] William Stephenson. Postulates of behaviorism. Philosophy of Science, 20(2):110–120, April 1953. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185447.

Sellars:1953:TSP

[209] Wilfrid Sellars. Is there a synthetic a Priori? Philosophy of Science, 20(2):121–138, April 1953. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185448.

Plochmann:1953:DTH

[210] George Kimball Plochmann. D’Arcy Thompson: His conception of the living body. Philosophy of Science, 20(2):139–148, April 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185449.

Herrmann:1953:ARB

[211] Heinz Herrmann. An account of recent biological methodology: Causal law and transplanar analysis. Philosophy of Science, 20(2):149–156, April 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185450.

Ballard:1953:RD

[212] Edward G. Ballard. The routine of discovery. Philosophy of Science, 20(2):157–163, April 1953. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185451.

Klein:1953:BRBa

[213] Martin J. Klein. Book review: The Nature of Some of Our Physical Concepts by P. W. Bridgman. Philosophy of Science, 20(2):164, April 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185454.

McGary:1953:BRBb

[214] Keith McGary. Book review: How to Develop Your Thinking Ability by Kenneth S. Keyes. Philosophy of Science, 20(2):164, April 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185452.

Portnoy:1953:BRB

[215] Julius Portnoy. Book review: An Introduction to Aesthetics by Hunter Mead. Philosophy of Science, 20(2):164, April 1953. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185453.

A:1953:BRBc

[216] R. L. A. Book review: A Critique of Logical Positivism by C. E. M. Joad. Philosophy of Science, 20(2):164–165, April 1953. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185455.

Litwak:1953:BRB

[217] Eleanor S. Litwak. Book review: The Hebrew Impact on Western Civilization by Dagobert D. Runes. Philosophy of Science, 20(2):165–166, April 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185456.

A:1953:BRBd

[218] R. L. A. Book review: The Perception of the Visual World by James J. Gibson. Philosophy of Science, 20(2):166, April 1953. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185457.

Golightly:1953:BRBb

[219] Cornelius L. Golightly. Book review: Two Commandments by Carl Jørgensen; Peter Hardt. Philosophy of Science, 20(2):166, April 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185458.

Anonymous:1953:FMb

[220] Anonymous. Front matter. Philosophy of Science, 20(2):??, April 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185444.

Benjamin:1953:STD

[221] A. Cornelius Benjamin. Some theories of the development of science. Philosophy of Science, 20(3):167–176, July 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185495.

Burks:1953:PTI

[222] Arthur W. Burks. The presupposition theory of induction. Philosophy of Science, 20(3):177–197, July 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185496.

Edel:1953:CVC

[223] Abraham Edel. Concept of values in contemporary philosophical value theory. Philosophy of Science, 20(3):198–207, July 1953. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185497.

Carmichael:1953:MMS

[224] Peter A. Carmichael. The metaphysical matrix of science. Philosophy of Science, 20(3):208–216, July 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185498.

Bar-Hillel:1953:SLP

[225] Yehoshua Bar-Hillel. Some linguistic problems connected with machine translation. Philosophy of Science, 20(3):217–225, July 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185499.

Conger:1953:CRW

[226] George P. Conger. Comments on a review by Werkmeister. Philosophy of Science, 20(3):226, July 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 185500. See [183].

McKinney:1953:CPP

[227] J. P. McKinney. Comment on a paper by Professor Kemble. Philosophy of Science, 20(3):227–231, July 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 185501. See [113, 228].

Kemble:1953:RMM

[228] E. C. Kemble. Reply to Mr. McKinney. Philosophy of Science, 20(3):232– 235, July 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185502. See [227].

Nordberg:1953:STT

[229] Robert Nordberg. A simple theory of time. Philosophy of Science, 20(3): 236–237, July 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185503.

Biser:1953:TE

[230] Erwin Biser. Time and events. Philosophy of Science, 20(3):238–240, July 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185504.

Anonymous:1953:NNa

[231] Anonymous. News and notices. Philosophy of Science, 20(3):241, July 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185505.

A:1953:BRBe

[232] R. L. A. Book review: Modern Science and Modern Man by James B. Conant. Philosophy of Science, 20(3):242, July 1953. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185506.

A:1953:BRBf

[233] R. L. A. Book review: Can We Agree? by Chauncey D. Leake; Patrick Romanell. Philosophy of Science, 20(3):242–243, July 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185507.

Leys:1953:BRB

[234] Wayne A. R. Leys. Book review: Butler’s Moral Philosophy by Austin Duncan-Jones. Philosophy of Science, 20(3):243, July 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185508.

Hall:1953:BRB

[235] Calvin S. Hall. Book review: Problems of Consciousness by H. A. Abramson. Philosophy of Science, 20(3):243–244, July 1953. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185509.

Anonymous:1953:BRa

[236] Anonymous. Books received. Philosophy of Science, 20(3):245, July 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185510.

Anonymous:1953:FMc

[237] Anonymous. Front matter. Philosophy of Science, 20(3):??, July 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185494.

Bittle:1953:LCA

[238] William E. Bittle. Language and culture areas: A note on method. Philosophy of Science, 20(4):247–256, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185029.

Churchman:1953:CP

[239] C. West Churchman. Concepts without primitives. Philosophy of Science, 20(4):257–265, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185030.

Mills:1953:TSR

[240] C. Wright Mills. Two styles of research in current social studies. Philosophy of Science, 20(4):266–275, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185031.

vanHeerden:1953:WM

[241] P. J. van Heerden. What is matter? Philosophy of Science, 20(4):276– 285, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185032.

Golightly:1953:JLT

[242] Cornelius L. Golightly. The James-lange theory: A logical Post-mortem. Philosophy of Science, 20(4):286–299, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/185033.

Chessick:1953:ANS

[243] Richard D. Chessick. The application of neurological studies in an approach to some philosophical problems. Philosophy of Science, 20(4):300– 312, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185034.

Hochberg:1953:EPR

[244] Herbert Hochberg. The empirical philosophy of roger and Francis bacon. Philosophy of Science, 20(4):313–326, October 1953. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185035.

Rescher:1953:MMG

[245] Nicholas Rescher. Mr. Madden on gestalt theory. Philosophy of Science, 20(4):327–328, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185036.

Madden:1953:SPG

[246] Edward H. Madden. Science, philosophy, and gestalt theory. Philosophy of Science, 20(4):329–331, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185037.

Silva:1953:OSS

[247] Herman Silva and George C. Vayonis. Objectivity and subjectivity in scientific research. Philosophy of Science, 20(4):332–334, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185038.

Lindley:1953:MND

[248] T. Foster Lindley. Moore’s nominal definitions of ’culture’. Philosophy of Science, 20(4):335–338, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185039.

Moore:1953:DLN

[249] Omar Khayyam Moore. Dr. Lindley and “nominal definitions of ‘culture’ ”. Philosophy of Science, 20(4):339–340, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185040.

Anonymous:1953:NNb

[250] Anonymous. News and notices. Philosophy of Science, 20(4):341, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185041.

Hoekstra:1953:BRB

[251] Raymond Hoekstra. Book review: Roger Bacon and His Search for a Universal Science by Stewart C. Easton. Philosophy of Science, 20(4):

342, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185042.

Linsky:1953:BRBb

[252] Leonard Linsky. Book review: The Vienna Circle: The Origin of Neo-Positivism by Victor Kraft. Philosophy of Science, 20(4):342, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185044.

Nassau:1953:BRB

[253] J. J. Nassau. Book review: The History of Astronomy by Giorgio Abetti; B. B. Abetti. Philosophy of Science, 20(4):342, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185043.

Linsky:1953:BRBc

[254] Leonard Linsky. Book review: Translations from the Philosophical Writings of Gottlob Frege by Peter Geach; Max Black. Philosophy of Science, 20(4):342–343, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185045.

Rudner:1953:BRB

[255] Richard Rudner. Book review: Semantics and the Philosophy of Language by Leonard Linsky. Philosophy of Science, 20(4):343, October 1953. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185046.

Wick:1953:BRB

[256] Warner Wick. Book review: The Moral Nature of Man by A. Campbell Garnett. Philosophy of Science, 20(4):343–344, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185047.

Klein:1953:BRBb

[257] Martin J. Klein. Book review: Quantum Theory of Matter by John C. Slater. Philosophy of Science, 20(4):344–345, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185048.

A:1953:BRBg

[258] R. L. A. Book review: Experimental Design in Psychological Research by Allen L. Edwards. Philosophy of Science, 20(4):345, October 1953.

CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185050.

A:1953:BRBh

[259] R. L. A. Book review: The Origins of Modern Science by Herbert Butterfield. Philosophy of Science, 20(4):345, October 1953. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185051.

Copi:1953:BRBb

[260] Irving Copi. Book review: Intentional Logic by Henry B. Veatch. Philosophy of Science, 20(4):345, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185049.

A:1953:BRBi

[261] R. L. A. Book review: Surveys, Polls, and Samples by Mildred B. Parten. Philosophy of Science, 20(4):345–346, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185052.

A:1953:BRBj

[262] R. L. A. Book review: Harvard Case Histories in Experimental Science by James B. Conant. Philosophy of Science, 20(4):346, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185053.

A:1953:BRBk

[263] R. L. A. Book review: The Development of Logical Empiricism by Joergen Joergensen. Philosophy of Science, 20(4):346, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185054.

A:1953:BRBl

[264] R. L. A. Book review: The Restless Universe by Max Born. Philosophy of Science, 20(4):346, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185055.

A:1953:BRBm

[265] R. L. A. Book review: Cybernetics: Circular Causal and Feedback Mechanisms in Biological and Social Systems by H. von Foerster. Philosophy of Science, 20(4):346–347, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185056.

A:1953:BRBn

[266] R. L. A. Book review: Mind, Perception and Science by W. Russell Brain. Philosophy of Science, 20(4):347, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185057.

A:1953:BRBo

[267] R. L. A. Book review: A Guide to the History of Science by George Sarton. Philosophy of Science, 20(4):347, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185058.

A:1953:BRBp

[268] R. L. A. Book review: Art and Technics by Lewis Mumford. Philosophy of Science, 20(4):347, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185059.

A:1953:BRBq

[269] R. L. A. Book reviews: Bertrand Russell’s Dictionary of Mind, Matter and Morals by Lester E. Denonn; Spinoza Dictionary by Dagobert D. Runes. Philosophy of Science, 20(4):347–348, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185060.

A:1953:BRBr

[270] R. L. A. Book review: The March toward Matter by John MacPartland. Philosophy of Science, 20(4):348, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185061.

A:1953:BRBs

[271] R. L. A. Book review: The Rational and the Superrational by Cassius J. Keyser. Philosophy of Science, 20(4):348, October 1953. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185062.

Anonymous:1953:BRb

[272] Anonymous. Books received. Philosophy of Science, 20(4):349, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185063.

Anonymous:1953:BM

[273] Anonymous. Back matter. Philosophy of Science, 20(4):350–358, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185064.

Anonymous:1953:VIb

[274] Anonymous. Volume information. Philosophy of Science, 20(4):359–360, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185027.

Anonymous:1953:FMd

[275] Anonymous. Front matter. Philosophy of Science, 20(4):??, October 1953. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185028.

Flower:1954:CET

[276] Elizabeth Flower. Comments on the ethical theory of Edgar A. Singer. Philosophy of Science, 21(1):1–8, January 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185855.

Kaminsky:1954:CL

[277] Jack Kaminsky. Corrigibility and law. Philosophy of Science, 21(1):9–15, January 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185856.

Axinn:1954:TCO

[278] Sidney Axinn. Two concepts of optimism. Philosophy of Science, 21(1): 16–24, January 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185857.

Brown:1954:ELS

[279] Robert Brown. Explanation by laws in social science. Philosophy of Science, 21(1):25–32, January 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185858.

Harris:1954:SCS

[280] Catherine Harris. Sullivan’s concept of scientific method as applied to psychiatry. Philosophy of Science, 21(1):33–43, January 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185859.

McNaughton:1954:ASC

[281] Robert McNaughton. Axiomatic systems, conceptual schemes, and the consistency of mathematical theories. Philosophy of Science, 21(1):44–53, January 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185860.

Haworth:1954:CVS

[282] L. Haworth and J. S. Minas. Concerning value science. Philosophy of Science, 21(1):54–61, January 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185861.

Plochmann:1954:QPQ

[283] George Kimball Plochmann. Is quantity prior to quality? Philosophy of Science, 21(1):62–67, January 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 185862. See comments [284].

Storer:1954:CPP

[284] Thomas Storer. Comments on Professor Plochmann’s “Is Quantity prior to Quality?”. Philosophy of Science, 21(1):68–73, January 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185863. See [283].

Anonymous:1954:NNa

[285] Anonymous. News and notices. Philosophy of Science, 21(1):74, January 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185864.

Nassau:1954:BRB

[286] J. J. Nassau. Book review: A History of Astronomy from Thales to Kepler by J. L. E. Dryer. Philosophy of Science, 21(1):75, January 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185867.

Portnoy:1954:BRBa

[287] Julius Portnoy. Book review: Vision and Action by Sidney Ratner. Philosophy of Science, 21(1):75, January 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185865.

VanVoorhis:1954:BRI

[288] W. R. Van Voorhis. Book review: An Introduction to Mathematical Thought by E. R. Stabler. Philosophy of Science, 21(1):75, January 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185866.

A:1954:BRBa

[289] R. L. A. Book review: The Scientific Adventure by Herbert Dingle. Philosophy of Science, 21(1):76, January 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185868.

A:1954:BRBb

[290] R. L. A. Book review: Shame and Guilt by Gerhart Piers; Milton B. Singer. Philosophy of Science, 21(1):76, January 1954. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185869.

A:1954:BRBc

[291] R. L. A. Book review: Evaluating Research and Development by I. R. Weschler; Paula Brown. Philosophy of Science, 21(1):76, January 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185870.

Anonymous:1954:BRa

[292] Anonymous. Books received. Philosophy of Science, 21(1):77, January 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185871.

Anonymous:1954:BMa

[293] Anonymous. Back matter. Philosophy of Science, 21(1):??, January 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185872.

Anonymous:1954:FMa

[294] Anonymous. Front matter. Philosophy of Science, 21(1):??, January 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185854.

Margenau:1954:CTF

[295] Henry Margenau. Can time flow backwards? Philosophy of Science, 21(2): 79–92, April 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185789.

Littauer:1954:SAS

[296] Sebastian B. Littauer. Social aspects of scientific method in industrial production. Philosophy of Science, 21(2):93–100, April 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185790.

Ushenko:1954:TSP

[297] A. P. Ushenko. Truth in science and in philosophy. Philosophy of Science, 21(2):101–117, April 1954. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185791.

Hartung:1954:CRM

[298] Frank E. Hartung. Cultural relativity and moral judgments. Philosophy of Science, 21(2):118–126, April 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185792.

Kogan:1954:MFN

[299] Zuce Kogan. Methods of furthering new ideas. Philosophy of Science, 21(2):127–131, April 1954. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185793.

Colburn:1954:LPR

[300] N. H. Colburn. Logic and professor ryle. Philosophy of Science, 21(2):132– 139, April 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185794.

Brodbeck:1954:PSS

[301] May Brodbeck. On the philosophy of the social sciences. Philosophy of Science, 21(2):140–156, April 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185795.

Gewirth:1954:SOS

[302] Alan Gewirth. Subjectivism and objectivism in the social sciences. Philosophy of Science, 21(2):157–163, April 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185796.

Rudner:1954:PSS

[303] Richard S. Rudner. Philosophy and social science. Philosophy of Science, 21(2):164–168, April 1954. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185797.

Nakhnikian:1954:NWD

[304] George Nakhnikian. A note on W. Donald Oliver’s theory of order. Philosophy of Science, 21(2):169–172, April 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185798.

Anonymous:1954:NNb

[305] Anonymous. News and notices. Philosophy of Science, 21(2):173, April 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185799.

Anonymous:1954:BRb

[306] Anonymous. Books received. Philosophy of Science, 21(2):174, April 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185800.

Anonymous:1954:BMb

[307] Anonymous. Back matter. Philosophy of Science, 21(2):??, April 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185801.

Anonymous:1954:FMb

[308] Anonymous. Front matter. Philosophy of Science, 21(2):??, April 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185788.

Singer:1954:DS

[309] E. A. Singer, Jr. Dialectic of the schools (i). Philosophy of Science, 21 (3):175–192, July 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185275.

King:1954:WD

[310] Lester S. King. What is disease? Philosophy of Science, 21(3):193– 203, July 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185276.

Sellars:1954:SRL

[311] Wilfrid Sellars. Some reflections on language games. Philosophy of Science, 21(3):204–228, July 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185277.

Gewirth:1954:CMC

[312] Alan Gewirth. Can Men change laws of social science? Philosophy of Science, 21(3):229–241, July 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185278.

Suppes:1954:SRP

[313] Patrick Suppes. Some remarks on problems and methods in the philosophy of science. Philosophy of Science, 21(3):242–248, July 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185279.

Grunbaum:1954:CPS

[314] Adolf Grünbaum. The clock paradox in the Special Theory of Relativity. Philosophy of Science, 21(3):249–253, July 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/ stable/185280. See comments [359, 404] and reply [405].

Pakswer:1954:IEI

[315] S. Pakswer. Information, entropy and inductive logic. Philosophy of Science, 21(3):254–259, July 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185281.

Lindley:1954:CFS

[316] T. Foster Lindley. The control factor in social experimentation. Philosophy of Science, 21(3):260–268, July 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185282.

Fries:1954:BRB

[317] Horace S. Fries. Book review: Methods of Inquiry: An Introduction to Philosophy and Scientific Method by C. West Churchman; Russell L. Ackoff. Philosophy of Science, 21(3):269, July 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185284.

Hare:1954:BRB

[318] Van Court Hare, Jr. Book review: The Mind of Kierkegaard by James Collins. Philosophy of Science, 21(3):269, July 1954. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185283.

Saunders:1954:BRB

[319] T. Frank Saunders. Book review: Conventional Logic and Modern Logic: A Prelude to Transition by Joseph T. Clark. Philosophy of Science, 21(3): 269, July 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185285.

Ruby:1954:BRB

[320] Lionel Ruby. Book review: Logic for Living by Henry Horace Williams. Philosophy of Science, 21(3):269–270, July 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185286.

A:1954:BRBd

[321] R. L. A. Book review: Hume: Theory of Knowledge by D. C. Yalden-Thomson. Philosophy of Science, 21(3):270, July 1954. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185290.

Glockler:1954:BRB

[322] George Glockler. Book review: A Source Book in Chemistry, 1400–1900 by H. M. Leicester; H. S. Klickstein. Philosophy of Science, 21(3):270, July 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185287.

Klein:1954:BRBa

[323] Martin J. Klein. Book review: Nuclear Physics by W. Heisenberg. Philosophy of Science, 21(3):270, July 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185289.

Portnoy:1954:BRBb

[324] Julius Portnoy. Book review: Philosophy and the Ideological Conflict by Charles S. Seely. Philosophy of Science, 21(3):270, July 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185288.

A:1954:BRBe

[325] R. L. A. Book review: Hume: Theory of Politics by Frederick Watkins. Philosophy of Science, 21(3):271, July 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185291.

A:1954:BRBf

[326] R. L. A. Book review: Berkeley: Philosophical Writings by T. E. Jessop. Philosophy of Science, 21(3):271, July 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185292.

A:1954:BRBg

[327] R. L. A. Book review: Alfred North Whitehead by F. S. C. Northrop; Mason Gross. Philosophy of Science, 21(3):271, July 1954. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185293.

A:1954:BRBh

[328] R. L. A. Book review: Readings in the Philosophy of Science by Philip P. Wiener. Philosophy of Science, 21(3):271, July 1954. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185294.

Kadish:1954:BRBa

[329] Mortimer R. Kadish. Book review: Philosophy and Psychoanalysis by John Wisdom. Philosophy of Science, 21(3):271, July 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185296.

Klein:1954:BRBb

[330] Martin J. Klein. Book review: Science in Progress (Seventh Series) by G. A. Baitsell. Philosophy of Science, 21(3):271, July 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185295.

Kadish:1954:BRBb

[331] Mortimer R. Kadish. Book review: Reason and Nature. An Essay on the Meaning of Scientific Method by Morris Raphael Cohen. Philosophy of Science, 21(3):271–272, July 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185297.

Eddison:1954:BRB

[332] R. T. Eddison. Book review: Genetics and the Origin of the Species by Theodosius Dobzhansky. Philosophy of Science, 21(3):272, July 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185298.

Anonymous:1954:BRc

[333] Anonymous. Books received. Philosophy of Science, 21(3):273, July 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185299.

Anonymous:1954:BMc

[334] Anonymous. Back matter. Philosophy of Science, 21(3):??, July 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185300.

Anonymous:1954:FMc

[335] Anonymous. Front matter. Philosophy of Science, 21(3):??, July 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185274.

Topitsch:1954:STP

[336] Ernst Topitsch. Society, technology, and philosophical reasoning. Philosophy of Science, 21(4):275–296, October 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/ stable/185951. See comment [537].

Singer:1954:DSI

[337] E. A. Singer, Jr. Dialectic of the schools (II). Philosophy of Science, 21 (4):297–315, October 1954. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185952.

Cartwright:1954:OTM

[338] Richard L. Cartwright. Ontology and the theory of meaning. Philosophy of Science, 21(4):316–325, October 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185953.

Bates:1954:MSD

[339] James Bates. A model for the science of decision. Philosophy of Science, 21(4):326–339, October 1954. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185954.

Simon:1954:ACM

[340] Herbert A. Simon. The axiomatization of classical mechanics. Philosophy of Science, 21(4):340–343, October 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185955.

Hamburg:1954:CNC

[341] Carl H. Hamburg. Critical note on “culture” and “learning-theory”. Philosophy of Science, 21(4):344–347, October 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185956.

Johnstone:1954:ASR

[342] Henry W. Johnstone, Jr. An alternative set of rules for the syllogism. Philosophy of Science, 21(4):348–351, October 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185957.

Anonymous:1954:NNc

[343] Anonymous. News and notices. Philosophy of Science, 21(4):352, October 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185958.

Nakhnikian:1954:BRB

[344] George Nakhnikian. Book review: Philosophical Investigations by Ludwig Wittgenstein; G. E. M. Anscombe; R. Rhees. Philosophy of Science, 21 (4):353–354, October 1954. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185959.

Ackoff:1954:BRBa

[345] Russel L. Ackoff. Book review: Great Tradition in Ethics by Ethel M. Albert; Theodore C. Denise; Sheldon P. Peterfreund. Philosophy of Science, 21(4):354, October 1954. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185960.

Ackoff:1954:BRBb

[346] Russell L. Ackoff. Book review: Main Currents of Scientific Thought: A History of the Sciences by Stephen S. Mason. Philosophy of Science, 21(4):354, October 1954. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185961.

Ackoff:1954:BRBc

[347] Russel L. Ackoff. Book review: An Introduction to Scientific Research by E. Bright Wilson. Philosophy of Science, 21(4):354, October 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185962.

Klein:1954:BRBc

[348] Martin J. Klein. Book review: The Aim and Structure of Physical Theory by Pierre Duhem; P. P. Wiener. Philosophy of Science, 21(4):354–355, October 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185963.

Anonymous:1954:BRd

[349] Anonymous. Books received. Philosophy of Science, 21(4):355–356, October 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/185964.

Anonymous:1954:VI

[350] Anonymous. Volume information. Philosophy of Science, 21(4):357–358, October 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185949.

Anonymous:1954:BMd

[351] Anonymous. Back matter. Philosophy of Science, 21(4):??, October 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185965.

Anonymous:1954:FMd

[352] Anonymous. Front matter. Philosophy of Science, 21(4):??, October 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185950.

Blum:1955:ARS

[353] Fred H. Blum. Action research — a scientific approach? Philosophy of Science, 22(1):1–7, January 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185762.

Barbour:1955:IFR

[354] Ian Barbour. Indeterminacy and freedom: A reappraisal. Philosophy of Science, 22(1):8–20, January 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185763.

ORourke:1955:ASA

[355] Neil W. O’Rourke. Amplifying systems and available energy. Philosophy of Science, 22(1):21–26, January 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185764.

Kemeny:1955:SP

[356] John G. Kemeny and Paul Oppenheim. Systematic power. Philosophy of Science, 22(1):27–33, January 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185765.

Gluck:1955:DSL

[357] Samuel E. Gluck. Do statistical laws have explanatory efficacy? Philosophy of Science, 22(1):34–38, January 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185766.

Scheffler:1955:SID

[358] Israel Scheffler. On synonymy and indirect discourse. Philosophy of Science, 22(1):39–44, January 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185767.

Leaf:1955:CPS

[359] Boris Leaf. The clock paradox in the Special Theory of Relativity. Philosophy of Science, 22(1):45–52, January 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/ stable/185768. See [314, 404, 405].

Grunbaum:1955:RDL

[360] Adolf Grünbaum. Reply to Dr. Leaf. Philosophy of Science, 22(1):53, January 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185769.

Koivisto:1955:MJV

[361] W. A. Koivisto. Moral judgments and value conflict. Philosophy of Science, 22(1):54–57, January 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185770.

Watkins:1955:MIR

[362] J. W. N. Watkins. Methodological individualism: A reply. Philosophy of Science, 22(1):58–62, January 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185771.

Benjamin:1955:BRB

[363] A. Cornelius Benjamin. Book review: Scientific Explanation: A Study of the Function of Theory, Probability and Law in Science by R. B. Braithwaite. Philosophy of Science, 22(1):63–65, January 1955. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185772.

McKinney:1955:BRB

[364] John C. McKinney. Book review: The Design of Social Research by Russell L. Ackoff. Philosophy of Science, 22(1):65, January 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185773.

Haworth:1955:BRBa

[365] Lawrence Haworth. Book reviews: Values and Policy in American Society by Russell E. Bayliff; Eugene Clark; Loyd Easton; Blaine E. Grimes; David H. Jennings; Norman H. Leonard; Readings in Social Policy by Bayliff; Problems in Social Policy by Bayliff. Philosophy of Science, 22(1):66, January 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185774.

Anonymous:1955:BRB

[366] Anonymous. Book review: Anton Lavoisier: Scientist, Economist, Social Reformer by Douglas McKie. Philosophy of Science, 22(1):66–67, January 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185775.

Ackoff:1955:BRBa

[367] Russell L. Ackoff. Book review: Theory and Method in the Social Sciences by Arnold M. Rose. Philosophy of Science, 22(1):67, January 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185776.

Ackoff:1955:BRBb

[368] Russell L. Ackoff. Book review: The Study of Behavior: Q-Technique and Its Methodology by William Stephenson. Philosophy of Science, 22(1):67, January 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185777.

Ackoff:1955:BRBc

[369] Russell L. Ackoff. Book review: Cybernetics (Transactions of the Ninth Conference, March 20-21, 1952) by H. Von Foerster. Philosophy of Science, 22(1):68, January 1955. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185779.

Hall:1955:BRBa

[370] Calvin S. Hall. Book review: Problems of Consciousness: Transactions of the Third Conference by H. A. Abramson. Philosophy of Science, 22(1): 68, January 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185778.

Hare:1955:BRBa

[371] Van Court Hare, Jr. Book review: Astrology and Alchemy by Mark Graubard. Philosophy of Science, 22(1):68, January 1955. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185781.

Klein:1955:BRB

[372] Martin J. Klein. Book review: Physics: Principles and Applications. Second Edition by Henry Margenau; William W. Watson; C. G. Montgomery. Philosophy of Science, 22(1):68, January 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185780.

Hare:1955:BRBb

[373] Van Court Hare, Jr. Book review: The Study of Human Nature by David Lindsay Watson. Philosophy of Science, 22(1):68–69, January 1955. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185782.

Hall:1955:BRBb

[374] Calvin S. Hall. Book review: Oneirics and Psychosomatics by Rolf Loehrich. Philosophy of Science, 22(1):69, January 1955. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185784.

Hare:1955:BRBc

[375] Van Court Hare, Jr. Book review: Sociological Abstracts by Leo P. Chall. Philosophy of Science, 22(1):69, January 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185783.

Anonymous:1955:BRa

[376] Anonymous. Books received. Philosophy of Science, 22(1):69–70, January 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185785.

Anonymous:1955:AR

[377] Anonymous. Abstracts received. Philosophy of Science, 22(1):71–72, January 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/185786.

Anonymous:1955:BMa

[378] Anonymous. Back matter. Philosophy of Science, 22(1):??, January 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185787.

Anonymous:1955:FMa

[379] Anonymous. Front matter. Philosophy of Science, 22(1):??, January 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185761.

Buck:1955:CJC

[380] Roger C. Buck and W. Seeman. Clinical judges and clinical insight in psychology. Philosophy of Science, 22(2):73–85, April 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185544.

Bar-Hillel:1955:EIT

[381] Yehoshua Bar-Hillel. An examination of information theory. Philosophy of Science, 22(2):86–105, April 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185545.

Sjoberg:1955:CMS

[382] Gideon Sjoberg. The comparative method in the social sciences. Philosophy of Science, 22(2):106–117, April 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/ stable/185546. See comments [383].

Rapoport:1955:CBC

[383] Anatol Rapoport. Comments on “The Comparative Method in the Social Sciences”. Philosophy of Science, 22(2):118–122, April 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185547. See [382].

Bock:1955:DST

[384] Kenneth E. Bock. Darwin and social theory. Philosophy of Science, 22(2): 123–134, April 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185548.

Frazer:1955:SIU

[385] William R. Frazer. Some indications of unity among the sciences. Philosophy of Science, 22(2):135–139, April 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185549.

Davidson:1955:OFT

[386] Donald Davidson, J. C. C. McKinsey, and Patrick Suppes. Outlines of a formal theory of value, I. Philosophy of Science, 22(2):140–160, April 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185550.

Denbow:1955:MFD

[387] Carl H. Denbow. Is mathematics a formal discipline? Philosophy of Science, 22(2):161–164, April 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185551.

Anonymous:1955:NNa

[388] Anonymous. News and notices. Philosophy of Science, 22(2):165, April 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185552.

Hinshaw:1955:BRB

[389] Virgil Hinshaw, Jr. Book review: The Metaphysics of Logical Positivism by Gustav Bergmann. Philosophy of Science, 22(2):166–167, April 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185553.

Martin:1955:BRB

[390] Norman M. Martin. Book review: Einführung in Die Symbolische Logik Mit Besonderer Berücksichtigung Ihrer Anwendungen by Rudolf Carnap. Philosophy of Science, 22(2):167–168, April 1955. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185554.

Geiger:1955:BRB

[391] George R. Geiger. Book reviews: Essays in Experimental Logic by John Dewey; Inquiry into Inquiries by Arthur F. Bentley; Sidney Ratner. Philosophy of Science, 22(2):168–171, April 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185555.

Kretzmann:1955:BRB

[392] Norman Kretzmann. Book review: A Budget of Paradoxes by Augustus De Morgan; David Eugene Smith. Philosophy of Science, 22(2):171, April 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185556.

Haworth:1955:BRBb

[393] Lawrence Haworth. Book review: Contemporary Philosophy by James Jarrett; Sterling McMurrin. Philosophy of Science, 22(2):172, April 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185558.

Portnoy:1955:BRB

[394] Julius Portnoy. Book review: The Compass of Philosophy by Newton P. Stallknecht; Robert S. Brumbaugh. Philosophy of Science, 22(2):172, April 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/185557.

Anonymous:1955:BRc

[395] Anonymous. Books received. Philosophy of Science, 22(2):173, April 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185559.

Anonymous:1955:BMb

[396] Anonymous. Back matter. Philosophy of Science, 22(2):??, April 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185560.

Anonymous:1955:FMb

[397] Anonymous. Front matter. Philosophy of Science, 22(2):??, April 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185543.

Burks:1955:DS

[398] Arthur W. Burks. Dispositional statements. Philosophy of Science, 22(3): 175–193, July 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185313.

Khatchadourian:1955:SMP

[399] Haig Khatchadourian. Some metaphysical presuppositions of science. Philosophy of Science, 22(3):194–204, July 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185314.

Ballard:1955:NUD

[400] Edward G. Ballard. On the nature and use of dialectic. Philosophy of Science, 22(3):205–213, July 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185315.

Salmon:1955:SR

[401] Wesley C. Salmon. The short run. Philosophy of Science, 22(3):214– 221, July 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185316. See comment

[454].

Seaman:1955:WR

[402] Francis Seaman. Whitehead and Relativity. Philosophy of Science, 22(3): 222–226, July 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185317.

Simon:1955:PHC

[403] Herbert A. Simon. Prediction and hindsight as confirmatory evidence. Philosophy of Science, 22(3):227–230, July 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185318.

Tornebohm:1955:CPG

[404] Håkan Törnebohm. Comments on Professor Grünbaum’s: “The Clock Paradox in the Special Theory of Relativity”. Philosophy of Science, 22 (3):231–232, July 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185319. See

[314, 359, 405].

Grunbaum:1955:RDT

[405] Adolf Grünbaum. Reply to Dr. Törnebohm’s comments on my article. Philosophy of Science, 22(3):233, July 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185320.

Nakhnikian:1955:BRB

[406] George Nakhnikian. Book review: Henri Poincaré: Critic of Crisis by Tobias Dantzig. Philosophy of Science, 22(3):234–235, July 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185321.

Suits:1955:BRB

[407] Bernard Suits. Book review: Aesthetics and Language by W. B. Gallie; Gilbert Ryle; Beryl Lake; Arnold Isenberg; Stuart Hampshire; J. A. Passmore; O. K. Bouwsma; Margaret McDonald; Helen Knight; Paul Ziff; William Elton. Philosophy of Science, 22(3):235–236, July 1955. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185322.

Harris:1955:BRBa

[408] H. S. Harris. Book review: The Mind and the Eye by Agnes Arber. Philosophy of Science, 22(3):236, July 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185323.

Stannard:1955:BRB

[409] Jerry Stannard. Book review: A Dictionary of Linguistics by Mario A. Pei; Frank Gaynor. Philosophy of Science, 22(3):236–237, July 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185324.

Harris:1955:BRBb

[410] Ruth Koski Harris. Book review: Adrenal Cortex by Elaine P. Ralli. Philosophy of Science, 22(3):237, July 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185326.

Haworth:1955:BRBc

[411] Lawrence Haworth. Book review: Historical Aspects of Organic Evolution by Philip G. Fothergill. Philosophy of Science, 22(3):237, July 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185325.

Feldstein:1955:BRBa

[412] Leonard C. Feldstein. Book review: Contributions to the Analysis and Synthesis of Knowledge by Else Frenkel-Brunswick. Philosophy of Science, 22(3):237–238, July 1955. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185327.

Feldstein:1955:BRBb

[413] Leonard C. Feldstein. Book review: The Origins and History of Consciousness by Erich Neumann; R. F. C. Hull. Philosophy of Science, 22 (3):238, July 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185328.

Doney:1955:BRB

[414] Willis Doney. Book review: Problems of Analysis by Max Black. Philosophy of Science, 22(3):238–241, July 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185329.

Anonymous:1955:BRd

[415] Anonymous. Books received. Philosophy of Science, 22(3):241–242, July 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185330.

Anonymous:1955:BMc

[416] Anonymous. Back matter. Philosophy of Science, 22(3):??, July 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185331.

Anonymous:1955:FMc

[417] Anonymous. Front matter. Philosophy of Science, 22(3):??, July 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185312.

vonBertalanffy:1955:ERC

[418] L. von Bertalanffy. An essay on the relativity of categories. Philosophy of Science, 22(4):243–263, October 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 185344. See comment [552].

McKinney:1955:GHM

[419] John C. McKinney. George H. Mead and the philosophy of science. Philosophy of Science, 22(4):264–271, October 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185345.

Goudge:1955:WP

[420] T. A. Goudge. What is a population? Philosophy of Science, 22(4):272– 279, October 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185346.

Oppenheim:1955:CU

[421] Felix E. Oppenheim. Control and unfreedom. Philosophy of Science, 22 (4):280–288, October 1955. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185347.

Bucklew:1955:STP

[422] John Bucklew. The subjective tradition in phenomenological psychology. Philosophy of Science, 22(4):289–299, October 1955. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185348.

Rieser:1955:RBI

[423] Max Rieser. Report on the “International Congress for the Philosophy of Science” in Zurich, Switzerland, August 23–28, 1954. Philosophy of Science, 22(4):300–308, October 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185349.

Cushen:1955:WGO

[424] Walter E. Cushen. War games and operations research. Philosophy of Science, 22(4):309–320, October 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185350.

Leaf:1955:VCV

[425] Boris Leaf. Vectorial composition of velocities in Relativity. Philosophy of Science, 22(4):321–324, October 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185351.

Cohn:1955:BRB

[426] Thomas S. Cohn. Book review: The Handbook of Social Psychology by Gardner Lindzey. Philosophy of Science, 22(4):325–326, October 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185352.

Hochberg:1955:BRB

[427] Herbert Hochberg. Book review: An Introduction to Deductive Logic by Hugues Leblanc. Philosophy of Science, 22(4):326–327, October 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185353.

Harris:1955:BRBc

[428] H. S. Harris. Book review: Conflict and The Web of Group-Affiliations by Georg Simmel; Kurt H. Wolff; Reinhard Bendix. Philosophy of Science, 22(4):327, October 1955. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185354.

Harris:1955:BRBd

[429] H. S. Harris. Book review: Leibniz by Ruth Lydia Saw. Philosophy of Science, 22(4):327–328, October 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185355.

Harris:1955:BRBe

[430] H. S. Harris. Book review: The Platonic Renaissance in England by Ernst Cassirer; James P. Pettegrove. Philosophy of Science, 22(4):328, October 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185356.

Harris:1955:BRBf

[431] H. S. Harris. Book review: The Sophists by Mario Untersteiner; Kathleen Freeman. Philosophy of Science, 22(4):328–329, October 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185357.

Anonymous:1955:BRe

[432] Anonymous. Books received. Philosophy of Science, 22(4):329–330, October 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/185358.

Anonymous:1955:NNb

[433] Anonymous. News and notices. Philosophy of Science, 22(4):331–332, October 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185359.

Anonymous:1955:VI

[434] Anonymous. Volume information. Philosophy of Science, 22(4):333–334, October 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185342.

Anonymous:1955:BMd

[435] Anonymous. Back matter. Philosophy of Science, 22(4):??, October 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185360.

Anonymous:1955:FMd

[436] Anonymous. Front matter. Philosophy of Science, 22(4):??, October 1955. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185343.

Boulding:1956:SCE

[437] Kenneth E. Boulding. Some contributions of economics to the general theory of value. Philosophy of Science, 23(1):1–14, January 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184993.

Danto:1956:EH

[438] Arthur C. Danto. On explanations in history. Philosophy of Science, 23 (1):15–30, January 1956. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184994.

Rothstein:1956:ILP

[439] Jerome Rothstein. Information, logic, and physics. Philosophy of Science, 23(1):31–35, January 1956. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184995.

Turner:1956:MLD

[440] Joseph Turner. Maxwell on the logic of dynamical explanation. Philosophy of Science, 23(1):36–47, January 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184996.

Shepard:1956:BRS

[441] Herbert A. Shepard. Basic research and the social system of pure science. Philosophy of Science, 23(1):48–57, January 1956. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184997.

Putnam:1956:DDC

[442] Hilary Putnam. A definition of degree of confirmation for very Rich languages. Philosophy of Science, 23(1):58–62, January 1956. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184998.

Weltfish:1956:PFR

[443] Gene Weltfish. The perspective for fundamental research in anthropology. Philosophy of Science, 23(1):63–73, January 1956. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184999.

Anonymous:1956:NNa

[444] Anonymous. News and notices. Philosophy of Science, 23(1):74, January 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185000.

Anonymous:1956:BMa

[445] Anonymous. Back matter. Philosophy of Science, 23(1):??, January 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185001.

Anonymous:1956:FMa

[446] Anonymous. Front matter. Philosophy of Science, 23(1):??, January 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184992.

Sellars:1956:PRR

[447] Roy Wood Sellars. Physical realism and Relativity: Unfinished business. Philosophy of Science, 23(2):75–81, April 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185593.

Hamburg:1956:PES

[448] Carl H. Hamburg. Psychology and the ethics of survival. Philosophy of Science, 23(2):82–89, April 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185594.

Sjoberg:1956:SCP

[449] Gideon Sjoberg. Science and changing publication patterns. Philosophy of Science, 23(2):90–96, April 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185595.

Freistadt:1956:DMF

[450] Hans Freistadt. Dialectical materialism: A friendly interpretation. Philosophy of Science, 23(2):97–110, April 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185596.

Palter:1956:PPS

[451] Robert Palter. Philosophic principles and scientific theory. Philosophy of Science, 23(2):111–135, April 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185597.

Kultgen:1956:BMS

[452] J. H. Kultgen. Boyle’s metaphysic of science. Philosophy of Science, 23(2):136–141, April 1956. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185598.

Hanson:1956:EPT

[453] Norwood Russell Hanson. On elementary particle theory. Philosophy of Science, 23(2):142–148, April 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185599.

Pettijohn:1956:SBS

[454] William C. Pettijohn. Salmon on “The Short Run”. Philosophy of Science, 23(2):149, April 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185600. See

[401].

Salmon:1956:RP

[455] Wesley C. Salmon. Reply to Pettijohn. Philosophy of Science, 23(2):150– 152, April 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185601.

Suppes:1956:NGC

[456] Patrick Suppes. Nelson Goodman on the concept of logical simplicity. Philosophy of Science, 23(2):153–159, April 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185602.

Rosen:1956:BRB

[457] Edward Rosen. Book review: Concepts of Space: The History of Theories of Space in Physics by Max Jammer. Philosophy of Science, 23(2):160– 162, April 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185603.

Hinshaw:1956:BRB

[458] Virgil Hinshaw, Jr. Book review: The Philosophy of Science by Pravas Jivan Chaudhury. Philosophy of Science, 23(2):162–164, April 1956. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185604.

Gluck:1956:BRB

[459] Samuel E. Gluck. Book review: Selected Papers in Statistics and Probability by Abraham Wald by T. W. Anderson. Philosophy of Science, 23(2):164–166, April 1956. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185605.

Shapiro:1956:BRB

[460] J. M. Shapiro. Book review: The Foundations of Statistics by Leonard J. Savage. Philosophy of Science, 23(2):166, April 1956. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185606.

Rado:1956:BRB

[461] Tibor Rado. Book review: Mathematics and Plausible Reasoning by G. Polya. Philosophy of Science, 23(2):167, April 1956. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185607.

Ralph:1956:BRB

[462] Philip Lee Ralph. Book review: The Judgment of History by Marie Collins Swabey. Philosophy of Science, 23(2):167–169, April 1956. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185608.

Stannard:1956:BRB

[463] Jerry Stannard. Book review: Epicurus and His Philosophy by Norman Wentworth De Witt. Philosophy of Science, 23(2):169, April 1956. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185609.

Golightly:1956:BRBa

[464] Cornelius L. Golightly. Book review: Ethical Judgment: The Use of Science in Ethics by Abraham Edel. Philosophy of Science, 23(2):169–170, April 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/185610.

Golightly:1956:BRBb

[465] Cornelius L. Golightly. Book review: The Emotive Theory of Ethics by Avrum Stroll. Philosophy of Science, 23(2):170, April 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185611.

Anonymous:1956:BR

[466] Anonymous. Books received. Philosophy of Science, 23(2):171–173, April 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185612.

Anonymous:1956:NNb

[467] Anonymous. News and notices. Philosophy of Science, 23(2):174, April 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185613.

Anonymous:1956:BMb

[468] Anonymous. Back matter. Philosophy of Science, 23(2):??, April 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185614.

Anonymous:1956:FMb

[469] Anonymous. Front matter. Philosophy of Science, 23(2):??, April 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185592.

Bergmann:1956:REL

[470] Gustav Bergmann. Russell’s examination of Leibniz examined. Philosophy of Science, 23(3):175–203, July 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185414.

Feibleman:1956:MAS

[471] James K. Feibleman. Mathematics and its applications in the sciences. Philosophy of Science, 23(3):204–215, July 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185415.

Moon:1956:EUT

[472] Parry Moon and Domina Eberle Spencer. On the establishment of a universal time. Philosophy of Science, 23(3):216–229, July 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185416. See remarks [506].

Lenz:1956:CDD

[473] John W. Lenz. Carnap on defining “degree of confirmation”. Philosophy of Science, 23(3):230–236, July 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185417.

Jeffrey:1956:VAS

[474] Richard C. Jeffrey. Valuation and acceptance of scientific hypotheses. Philosophy of Science, 23(3):237–246, July 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185418.

Churchman:1956:SDM

[475] C. West Churchman. Science and decision making. Philosophy of Science, 23(3):247–249, July 1956. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185419.

Hochberg:1956:OO

[476] Herbert Hochberg. The ontological operator. Philosophy of Science, 23 (3):250–259, July 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185420.

Cartwright:1956:CDH

[477] Richard L. Cartwright. Comments on Dr. Hochberg’s paper. Philosophy of Science, 23(3):260–265, July 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185421.

Dunayevskaya:1956:LE

[478] Raya Dunayevskaya. Letter to Editor. Philosophy of Science, 23(3):266, July 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/185422.

Nakhnikian:1956:BRB

[479] George Nakhnikian. Book reviews: Introduction to Logic by Irving M. Copi; Symbolic Logic by Irving M. Copi. Philosophy of Science, 23(3):

267–268, July 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185423.

Harris:1956:BRB

[480] H. S. Harris. Book review: Science and the Human Imagination by Mary B. Hesse. Philosophy of Science, 23(3):268–269, July 1956. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185424.

Frankel:1956:BRBa

[481] Charles Frankel. Book review: The Little Community: Viewpoints for the Study of a Human Whole by Robert Redfield. Philosophy of Science, 23(3):269–270, July 1956. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185425.

Frankel:1956:BRBb

[482] Charles Frankel. Book review: Social Norms and Roles by Ragnar Rommetveit. Philosophy of Science, 23(3):270, July 1956. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185426.

Kaelin:1956:BRB

[483] E. F. Kaelin. Book reviews: On the Knowledge of Good and Evil by Philip Blair Rice; The Value Judgement by W. D. Lamont. Philosophy of Science, 23(3):270–272, July 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185427.

Anonymous:1956:NNc

[484] Anonymous. News and notices. Philosophy of Science, 23(3):273, July 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185428.

Anonymous:1956:BMc

[485] Anonymous. Back matter. Philosophy of Science, 23(3):??, July 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185429.

Anonymous:1956:FMc

[486] Anonymous. Front matter. Philosophy of Science, 23(3):??, July 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185413.

Sellars:1956:GRA

[487] Roy Wood Sellars. Gestalt and Relativity: An analogy. Philosophy of Science, 23(4):275–279, October 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184975.

Reiser:1956:PEB

[488] Oliver L. Reiser. Postulates for an ethics of belief in science, religion, and philosophy. Philosophy of Science, 23(4):280–282, October 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184976.

Wyman:1956:WPS

[489] Mary A. Wyman. Whitehead’s philosophy of science in the light of Wordsworth’s poetry. Philosophy of Science, 23(4):283–296, October 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184977.

Palter:1956:OO

[490] Robert Palter. Operations and the occult. Philosophy of Science, 23 (4):297–314, October 1956. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184978.

Walcott:1956:IM

[491] Gregory Dexter Walcott. An inquiring mind. Philosophy of Science, 23 (4):315–324, October 1956. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184979.

Handy:1956:PFI

[492] Rollo Handy. Personality factors and intellectual production. Philosophy of Science, 23(4):325–332, October 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/ stable/184980. See comments [526].

Harrah:1956:MC

[493] David Harrah. A model of communication. Philosophy of Science, 23 (4):333–342, October 1956. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184981.

Roder:1956:AWV

[494] Ferdinand Roder. An analysis of Warburg’s view on the origin of cancer cells. Philosophy of Science, 23(4):343–347, October 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184982.

Anonymous:1956:NNd

[495] Anonymous. News and notices. Philosophy of Science, 23(4):348, October 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184983.

Anonymous:1956:VI

[496] Anonymous. Volume information. Philosophy of Science, 23(4):349–350, October 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184973.

Anonymous:1956:BMd

[497] Anonymous. Back matter. Philosophy of Science, 23(4):??, October 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184984.

Anonymous:1956:FMd

[498] Anonymous. Front matter. Philosophy of Science, 23(4):??, October 1956. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184974.

Finch:1957:VRP

[499] Henry Albert Finch. Validity rules for proportionally quantified syllogisms. Philosophy of Science, 24(1):1–18, January 1957. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184928.

Miller:1957:IPC

[500] David L. Miller. The importance of presents in contemporary science. Philosophy of Science, 24(1):19–24, January 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184929.

Freistadt:1957:DMF

[501] Hans Freistadt. Dialetical materialism: A further discussion. Philosophy of Science, 24(1):25–40, January 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184930.

Alston:1957:SDL

[502] William P. Alston. Is a sense-datum language necessary? Philosophy of Science, 24(1):41–45, January 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

184931.

Wein:1957:TPA

[503] Hermann Wein. Trends in philosophical anthropology and cultural anthropology in postwar Germany. Philosophy of Science, 24(1):46–56, January 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184932.

Tornebohm:1957:ERS

[504] Håkan Törnebohm. Epistemological reflexions over the Special Theory of Relativity and Milne’s conception of two times. Philosophy of Science, 24(1):57–69, January 1957. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184933.

Dewan:1957:OMA

[505] Edmond M. Dewan. “other minds”: An application of recent epistemological ideas to the definition of consciousness. Philosophy of Science, 24(1):70–76, January 1957. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184934.

Grunbaum:1957:RCM

[506] Adolf Grünbaum. Remarks concerning Moon and Spencer’s “On the Establishment of a Universal Time”. Philosophy of Science, 24(1):77–78, January 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184935. See [472].

Capek:1957:NAW

[507] Milic Capek. Note about Whitehead’s definitions of co-presence. Philosophy of Science, 24(1):79–86, January 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184936.

Bedau:1957:BRBa

[508] H. A. Bedau. Book review: Conceptual Thinking: A Logical Inquiry by Stephan Körner. Philosophy of Science, 24(1):87–89, January 1957. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184937.

Bromberger:1957:BRB

[509] Sylvain Bromberger. Book review: Operationism by A. Cornelius Benjamin. Philosophy of Science, 24(1):89–90, January 1957. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184938.

Golightly:1957:BRB

[510] Cornelius L. Golightly. Book review: The Logic of Moral Discourse by Paul Edwards. Philosophy of Science, 24(1):90–91, January 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184939.

Bedau:1957:BRBb

[511] H. A. Bedau. Book review: On Human Thinking by K. W. Monsarrat. Philosophy of Science, 24(1):91–92, January 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184940.

M:1957:BRB

[512] J. S. M. Book review: Meaning and Necessity: A Study in Semantics and Modal Logic by Rudolf Carnap. Philosophy of Science, 24(1):92, January 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184941.

Anonymous:1957:BRa

[513] Anonymous. Books received. Philosophy of Science, 24(1):93–94, January 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184942.

Anonymous:1957:NNa

[514] Anonymous. News and notices. Philosophy of Science, 24(1):95, January 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184943.

Anonymous:1957:BMa

[515] Anonymous. Back matter. Philosophy of Science, 24(1):??, January 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184944.

Anonymous:1957:FMa

[516] Anonymous. Front matter. Philosophy of Science, 24(1):??, January 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184927.

Spence:1957:EBT

[517] Kenneth W. Spence. The empirical basis and theoretical structure of psychology. Philosophy of Science, 24(2):97–108, April 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185715.

Feuer:1957:PS

[518] Lewis S. Feuer. The principle of simplicity. Philosophy of Science, 24(2): 109–122, April 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185716.

Pap:1957:BP

[519] Arthur Pap. Belief and propositions. Philosophy of Science, 24(2):123– 136, April 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185717.

Frankel:1957:EIH

[520] Charles Frankel. Explanation and interpretation in history. Philosophy of Science, 24(2):137–155, April 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185718.

Goldstein:1957:LEM

[521] Leon J. Goldstein. The logic of explanation in malinowskian anthropology. Philosophy of Science, 24(2):156–166, April 1957. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185719.

Madden:1957:ATP

[522] Edward H. Madden. Aristotle’s treatment of probability and signs. Philosophy of Science, 24(2):167–172, April 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185720.

Kaminsky:1957:CES

[523] Jack Kaminsky. Can “essence” be a scientific term? Philosophy of Science, 24(2):173–179, April 1957. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185721.

Salmon:1957:PI

[524] Wesley C. Salmon. The predictive inference. Philosophy of Science, 24(2): 180–190, April 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185722.

Hochberg:1957:PQP

[525] Herbert Hochberg. Professor Quine, Pegasus, and Dr. Cartwright. Philosophy of Science, 24(2):191–203, April 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185723.

Robinson:1957:CRH

[526] Bryan W. Robinson. Comments on Rollo Handy’s “personality factors and intellectual production”. Philosophy of Science, 24(2):204, April 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185724. See [492].

Anonymous:1957:ABJa

[527] Anonymous. Abstracts from the British journal for the philosophy of science. Philosophy of Science, 24(2):205, April 1957. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185725.

Anonymous:1957:BMb

[528] Anonymous. Back matter. Philosophy of Science, 24(2):??, April 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185726.

Anonymous:1957:FMb

[529] Anonymous. Front matter. Philosophy of Science, 24(2):??, April 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185714.

Hay:1957:FWP

[530] William H. Hay. Free-will and possibilities. Philosophy of Science, 24(3): 207–214, July 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185141.

Hiz:1957:TE

[531] H. Hiz. Types and environments. Philosophy of Science, 24(3):215–220, July 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/185142.

Caws:1957:RCS

[532] Peter Caws. A reappraisal of the conceptual Scheme of science. Philosophy of Science, 24(3):221–234, July 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185143.

Gellner:1957:ILK

[533] Ernest Gellner. Ideal language and kinship structure. Philosophy of Science, 24(3):235–242, July 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185144.

Patin:1957:PIF

[534] Henry A. Patin. Pragmatism, intuitionism, and formalism. Philosophy of Science, 24(3):243–252, July 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185145.

Delgado:1957:PMI

[535] Rafael Rodriguez Delgado. A possible model for ideas. Philosophy of Science, 24(3):253–269, July 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185146.

Reiser:1957:MAM

[536] Oliver L. Reiser. Matter, Anti-matter, and cosmic symmetry. Philosophy of Science, 24(3):271–274, July 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185147.

McKinney:1957:RRC

[537] J. P. McKinney. The rational and the real: Comment on a paper by E. Topitsch. Philosophy of Science, 24(3):275–280, July 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185148. See [336].

Biser:1957:BRBa

[538] Erwin Biser. Book review: Physics and Microphysics by Louis DeBroglie; Martin Davidson. Philosophy of Science, 24(3):281–282, July 1957. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185149.

Kaelin:1957:BRBa

[539] E. F. Kaelin. Book review: Ethical Value by George F. Hourani. Philosophy of Science, 24(3):282–284, July 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185150.

Parsons:1957:BRB

[540] Howard L. Parsons. Book review: Varieties of Human Value by Charles Morris. Philosophy of Science, 24(3):284–287, July 1957. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185151.

Gluck:1957:BRB

[541] Samuel E. Gluck. Book review: Industrial Society: The Emergence of the Human Problems of Automation by Georges Friedmann; Harold L. Sheppard. Philosophy of Science, 24(3):287–289, July 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185152.

Bedau:1957:BRBc

[542] Hugo A. Bedau. Book review: Language, Thought, and Reality: Selected Writings of Benjamin Lee Whorf by John B. Carroll. Philosophy of Science, 24(3):289–293, July 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185153.

Anonymous:1957:ABJb

[543] Anonymous. Abstracts from the British journal for the philosophy of science. Philosophy of Science, 24(3):294, July 1957. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185154.

Anonymous:1957:NNb

[544] Anonymous. News and notices. Philosophy of Science, 24(3):295, July 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185155.

Anonymous:1957:BMc

[545] Anonymous. Back matter. Philosophy of Science, 24(3):??, July 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185156.

Anonymous:1957:FMc

[546] Anonymous. Front matter. Philosophy of Science, 24(3):??, July 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185140.

Margenau:1957:REA

[547] Henry Margenau and Richard A. Mould. Relativity: An epistemological appraisal. Philosophy of Science, 24(4):297–307, October 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185004.

Lande:1957:NQF

[548] Alfred Landé. Non-quantal foundations of quantum theory. Philosophy of Science, 24(4):309–320, October 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185005.

McKnight:1957:QTC

[549] John L. McKnight. The quantum theoretical concept of measurement. Philosophy of Science, 24(4):321–330, October 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/185006.

Fitch:1957:CLW

[550] Frederic B. Fitch. Combinatory logic and Whitehead’s theory of prehensions. Philosophy of Science, 24(4):331–335, October 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185007.

Grunberg:1957:NVE

[551] Emile Grunberg. Notes on the verifiability of economic laws. Philosophy of Science, 24(4):337–348, October 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185008.

McKinney:1957:KEC

[552] J. P. McKinney. Knowledge and experience: Comment on a paper by L. von Bertalanffy on “The Relativity of Categories”. Philosophy of Science, 24(4):349–356, October 1957. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185009.

See [418].

Biser:1957:BRBb

[553] Erwin Biser. Book review: The Philosophy of Mathematics by Edward A. Maziarz. Philosophy of Science, 24(4):357, October 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185010.

Salmon:1957:BRBa

[554] Wesley C. Salmon. Book review: The Foundations of Science by Sheldon J. Lachman. Philosophy of Science, 24(4):358–359, October 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185011.

Biser:1957:BRBc

[555] Erwin Biser. Book review: The Philosophy of Science: An Introduction by Stephen Toulmin. Philosophy of Science, 24(4):359–360, October 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185012.

Kaelin:1957:BRBb

[556] E. F. Kaelin. Book review: Scientific Humanism and Christian Thought by D. Dubarle; Reginald Trevett. Philosophy of Science, 24(4):360–361, October 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185013.

Diesing:1957:BRB

[557] Paul Diesing. Book review: The Structure of a Moral Code by John Ladd. Philosophy of Science, 24(4):361–363, October 1957. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185014.

Podolsky:1957:BRB

[558] Boris Podolsky. Book review: Foundations of Quantum Theory — A Study in Continuity and Symmetry by Alfred Landé. Philosophy of Science, 24 (4):363–364, October 1957. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185015.

Salmon:1957:BRBb

[559] Wesley C. Salmon. Book review: Scientific Inference by Harold Jeffreys. Philosophy of Science, 24(4):364–366, October 1957. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185016.

Buchi:1957:BRB

[560] J. Richard Buchi. Book review: The Basic Concepts of Mathematics by Karl Menger. Philosophy of Science, 24(4):366, October 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185017.

Hinshaw:1957:BRB

[561] Virgil Hinshaw, Jr. Book review: The Revolution in Philosophy by A. J. Ayer; W. C. Kneale; G. A. Paul; D. F. Pears; P. F. Strawson; G. J. Warnock; R. A. Wollheim. Philosophy of Science, 24(4):366–367, October 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185018.

Moore:1957:BRB

[562] Edward C. Moore. Book review: The Leibniz-Clarke Correspondence by H. G. Alexander. Philosophy of Science, 24(4):367–369, October 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185019.

Kretzmann:1957:BRB

[563] Norman Kretzmann. Book review: Dominant Themes of Modern Philosophy. A History by George Boas. Philosophy of Science, 24(4):369–371, October 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185020.

Newmark:1957:BRB

[564] L. D. Newmark. Book review: The Origins and Prehistory of Language by G. Révész; J. Butler. Philosophy of Science, 24(4):371–372, October 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185021.

Anonymous:1957:BRb

[565] Anonymous. Books received. Philosophy of Science, 24(4):373–374, October 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/185022.

Anonymous:1957:CBP

[566] Anonymous. Correction: Belief and propositions. Philosophy of Science, 24(4):375, October 1957. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185024.

Anonymous:1957:NNc

[567] Anonymous. News and notices. Philosophy of Science, 24(4):375, October 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185023.

Anonymous:1957:ABJc

[568] Anonymous. Abstracts from the British journal for the philosophy of science. Philosophy of Science, 24(4):376, October 1957. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185025.

Anonymous:1957:VI

[569] Anonymous. Volume information. Philosophy of Science, 24(4):377–379, October 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185002.

Anonymous:1957:BMd

[570] Anonymous. Back matter. Philosophy of Science, 24(4):??, October 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185026.

Anonymous:1957:FMd

[571] Anonymous. Front matter. Philosophy of Science, 24(4):??, October 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185003.

Brodbeck:1958:MID

[572] May Brodbeck. Methodological individualisms: Definition and reduction. Philosophy of Science, 25(1):1–22, January 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185333.

Margenau:1958:PPC

[573] Henry Margenau. Philosophical problems concerning the meaning of measurement in physics. Philosophy of Science, 25(1):23–33, January 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185334.

Mikulak:1958:SPC

[574] Maxim W. Mikulak. Soviet philosophic-cosmological thought. Philosophy of Science, 25(1):35–50, January 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185335.

Hinshaw:1958:OH

[575] Virgil Hinshaw, Jr. The objectivity of history. Philosophy of Science, 25(1):51–58, January 1958. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185336.

Swartz:1958:HSA

[576] Marc J. Swartz. History and science in anthropology. Philosophy of Science, 25(1):59–70, January 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185337.

Myers:1958:PIP

[577] Charles M. Myers. Phenomenological idiom and perceptual mode. Philosophy of Science, 25(1):71–81, January 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185338.

Rescher:1958:TE

[578] Nicholas Rescher. A theory of evidence. Philosophy of Science, 25(1):83– 94, January 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185339.

Anonymous:1958:NNa

[579] Anonymous. News and notices. Philosophy of Science, 25(1):95–96, January 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/185340.

Anonymous:1958:BMa

[580] Anonymous. Back matter. Philosophy of Science, 25(1):??, January 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185341.

Anonymous:1958:FMa

[581] Anonymous. Front matter. Philosophy of Science, 25(1):??, January 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185332.

Hartman:1958:VFS

[582] Robert S. Hartman. Value, fact and science. Philosophy of Science, 25 (2):97–108, April 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185431.

Mayo:1958:IC

[583] Bernard Mayo. The incongruity of counterparts. Philosophy of Science, 25(2):109–115, April 1958. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185432.

Handy:1958:PNS

[584] Rollo Handy. Philosophy’s neglect of the social sciences. Philosophy of Science, 25(2):117–124, April 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185433.

Putman:1958:FCA

[585] Hilary Putman and Hilary Putnam. Formalization of the concept “about”. Philosophy of Science, 25(2):125–130, April 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/185434.

Apostle:1958:MSA

[586] H. G. Apostle. Methodological superiority of Aristotle over Euclid. Philosophy of Science, 25(2):131–134, April 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185435.

Hinshaw:1958:BRB

[587] Virgil Hinshaw, Jr. Book review: The Language of Modern Physics. An Introduction to the Philosophy of Science by Ernest H. Hutten. Philosophy of Science, 25(2):135–136, April 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185436.

Bedau:1958:BRBa

[588] Hugo A. Bedau. Book review: Logic and Knowledge: Essays 1901–1950 by Bertrand Russell; Robert Charles Marsh. Philosophy of Science, 25(2): 136–139, April 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185437.

Bedau:1958:BRBb

[589] Hugo Adam Bedau. Book review: The Psychology of Thought and Judgment by Donald M. Johnson. Philosophy of Science, 25(2):139–140, April 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185438.

Lucca:1958:BRBa

[590] John De Lucca. Book review: Life and Mind by Edmund Ware Sinnott. Philosophy of Science, 25(2):140, April 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185439.

Harris:1958:BRBa

[591] Ruth Koski Harris. Book review: Biochemical Individuality: The Basis for the Genetotrophic Concept by Roger J. Williams. Philosophy of Science, 25(2):140–141, April 1958. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185440.

Harris:1958:BRBb

[592] H. S. Harris. Book review: Whitehead’s Philosophical Development by Nathaniel Lawrence. Philosophy of Science, 25(2):141–142, April 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185441.

Lucca:1958:BRBb

[593] John De Lucca. Book review: Philosophic Problems by Maurice Mandelbaum; Francis W. Gramlich; Alan Ross Anderson. Philosophy of Science, 25(2):142, April 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185442.

Anonymous:1958:BMb

[594] Anonymous. Back matter. Philosophy of Science, 25(2):??, April 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185443.

Anonymous:1958:FMb

[595] Anonymous. Front matter. Philosophy of Science, 25(2):??, April 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185430.

Stopes-Roe:1958:SCC

[596] Harry V. Stopes-Roe. Some considerations concerning “interpretative systems”. Philosophy of Science, 25(3):143–156, July 1958. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185740.

Fontaine:1958:MER

[597] William T. Fontaine. The means end relation and its significance for crosscultural ethical agreement. Philosophy of Science, 25(3):157–162, July 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185741.

Senior:1958:VL

[598] James K. Senior. The vernacular of the laboratory. Philosophy of Science, 25(3):163–168, July 1958. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185742.

Bachem:1958:EEB

[599] A. Bachem. Ethics and esthetics on a biological basis. Philosophy of Science, 25(3):169–175, July 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185743.

Kultgen:1958:PCM

[600] J. H. Kultgen. Philosophic conceptions in Mendeleev’s Principles of Chemistry. Philosophy of Science, 25(3):177–183, July 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185744.

Kent:1958:SN

[601] William Kent. Scientific naming. Philosophy of Science, 25(3):185–193, July 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/185745.

Ito:1958:BEM

[602] Schuntaro Ito. Biologische erkenntnis und moderne physik. Philosophy of Science, 25(3):195–197, July 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185746.

Skarsgaard:1958:SRL

[603] Lars Skarsgård. Some remarks on the logic of explanation. Philosophy of Science, 25(3):199–207, July 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185747.

McKnight:1958:ELI

[604] John Lacy McKnight. An extended latency interpretation of quantum mechanical measurement. Philosophy of Science, 25(3):209–222, July 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185748.

Benjamin:1958:BRB

[605] A. Cornelius Benjamin. Book review: The Philosophy of Science by P. Henry van Laer; Henry J. Koren. Philosophy of Science, 25(3):223, July 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185749.

Jacobsen:1958:BRB

[606] T. S. Jacobsen. Book review: Space, Time, and Creation: Philosophical Aspects of Scientific Cosmology by Milton K. Munitz. Philosophy of Science, 25(3):223–226, July 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185750.

Gluck:1958:BRBa

[607] Samuel E. Gluck. Book review: Roots of Scientific Thought: A Cultural Perspective by P. P. Wiener; A. Noland. Philosophy of Science, 25(3):226– 228, July 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185751.

Haworth:1958:BRBa

[608] Lawrence L. Haworth. Book review: The Scientific Study of Social Behaviour by Michael Argyle. Philosophy of Science, 25(3):228–229, July 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185752.

Golightly:1958:BRBa

[609] Cornelius L. Golightly. Book review: Operations of Sociological Inquiry by Mario Lins. Philosophy of Science, 25(3):229, July 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185753.

Golightly:1958:BRBb

[610] Cornelius L. Golightly. Book review: Evolution: The Ages and Tomorrow by G. Murray McKinley. Philosophy of Science, 25(3):229–230, July 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185754.

Golightly:1958:BRBc

[611] Cornelius L. Golightly. Book review: Group Processes: Transactions of the Second Conference on Group Processes by Bertram Schaffner. Philosophy of Science, 25(3):230–231, July 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185755.

Gluck:1958:BRBb

[612] Samuel E. Gluck. Book review: System of Ethics by Leonard Nelson; Norbert Guterman. Philosophy of Science, 25(3):231–232, July 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185756.

Anonymous:1958:AD

[613] Anonymous. Abstracts from dialectica. Philosophy of Science, 25(3):233– 238, July 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185757.

Anonymous:1958:ABJ

[614] Anonymous. Abstracts from the British journal for the philosophy of science. Philosophy of Science, 25(3):238–239, July 1958. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185758.

Anonymous:1958:NNb

[615] Anonymous. News and notices. Philosophy of Science, 25(3):240, July 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185759.

Anonymous:1958:BMc

[616] Anonymous. Back matter. Philosophy of Science, 25(3):??, July 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185760.

Anonymous:1958:FMc

[617] Anonymous. Front matter. Philosophy of Science, 25(3):??, July 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185739.

Riepe:1958:FSN

[618] Dale Riepe. Flexible scientific naturalism and dialectical fundamentalism. Philosophy of Science, 25(4):241–248, October 1958. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185636.

Walsh:1958:SCE

[619] V. C. Walsh. Scarcity and the concepts of ethics. Philosophy of Science, 25(4):249–257, October 1958. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185637.

Wiedorn:1958:MRP

[620] Wm. S. Wiedorn, Jr. Method in research in psychiatry: Implications for the philosophy of science. Philosophy of Science, 25(4):259–262, October 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185638.

Fitch:1958:RSC

[621] Frederic B. Fitch. Representation of sequential circuits in combinatory logic. Philosophy of Science, 25(4):263–279, October 1958. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185639.

Copi:1958:BFP

[622] Irving M. Copi. The Burali-Forti Paradox. Philosophy of Science, 25 (4):281–286, October 1958. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185640.

Moon:1958:RC

[623] Parry Moon and Domina Eberle Spencer. Retardation in cosmology. Philosophy of Science, 25(4):287–292, October 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185641.

Rapoport:1958:BRB

[624] Anatol Rapoport. Book review: La Mathématique Sociale du Marquis de Condorcet by Gilles-Gaston Granger. Philosophy of Science, 25(4):293– 295, October 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185642.

Veatch:1958:BRB

[625] Henry Veatch. Book review: Logic without Metaphysics: And Other Essays in the Philosophy of Science by Ernest Nagel. Philosophy of Science, 25(4):295–297, October 1958. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185643.

Wiener:1958:BRB

[626] Philip P. Wiener. Book review: The Copernican Revolution: Planetary Astronomy in the Development of Western Thought by Thomas S. Kuhn. Philosophy of Science, 25(4):297–299, October 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185644.

Haworth:1958:BRBb

[627] Lawrence Haworth. Book review: A Natural Science of Society by A. R. Radcliffe-Brown. Philosophy of Science, 25(4):299–300, October 1958.

CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185645.

Stannard:1958:BRB

[628] Jerry Stannard. Book reviews: Milton’s Ontology, Cosmogony and Physics by Walter Clyde Curry; Milton and Science by Kester Svendsen. Philosophy of Science, 25(4):300–301, October 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185646.

Madden:1958:BRB

[629] Edward H. Madden. Book review: Science and the Creative Spirit by Karl W. Deutsch; F. E. L. Priestley; Harcourt Brown; David Hawkins; American Council of Learned Societies. Philosophy of Science, 25(4):301– 302, October 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185647.

Yolton:1958:BRB

[630] John W. Yolton. Book review: Perceiving: A Philosophical Study by Roderick M. Chisholm. Philosophy of Science, 25(4):302–305, October 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185648.

Johnstone:1958:BRB

[631] Henry W. Johnstone, Jr. Book review: British Philosophy in the Mid-Century by C. A. Mace. Philosophy of Science, 25(4):305–307, October 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185649.

Kaelin:1958:BRBa

[632] E. F. Kaelin. Book review: The Language of Value by Ray Lepley. Philosophy of Science, 25(4):307–308, October 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185650.

Dickie:1958:BRB

[633] George T. Dickie. Book review: The Age of Ideology by Henry D. Aiken. Philosophy of Science, 25(4):308–309, October 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185651.

Kaelin:1958:BRBb

[634] E. F. Kaelin. Book review: Physics and Metaphysics of Music, and Essays on the Philosophy of Mathematics by Lazare Saminsky. Philosophy of Science, 25(4):309, October 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185652.

Minas:1958:BRB

[635] J. S. Minas. Book review: Inquiry: An Interdisciplinary Journal of Philosophy and the Social Sciences. Volume 1, Number 1 by Arne Naess. Philosophy of Science, 25(4):309–310, October 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185653.

Anonymous:1958:VI

[636] Anonymous. Volume information. Philosophy of Science, 25(4):311–312, October 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185634.

Anonymous:1958:BMd

[637] Anonymous. Back matter. Philosophy of Science, 25(4):??, October 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185654.

Anonymous:1958:FMd

[638] Anonymous. Front matter. Philosophy of Science, 25(4):??, October 1958. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185635.

Williams:1959:EHN

[639] Thomas Rhys Williams. The evolution of a human nature. Philosophy of Science, 26(1):1–13, January 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185570.

Darlington:1959:CL

[640] Jared Darlington. On the confirmation of laws. Philosophy of Science, 26(1):14–24, January 1959. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185571.

See comments [707, 705].

Resnick:1959:CH

[641] Lawrence Resnick. Confirmation and hypothesis. Philosophy of Science, 26(1):25–30, January 1959. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185572.

Bork:1959:MES

[642] Alfred Morton Bork. Methodology of the empirical sciences. Philosophy of Science, 26(1):31–34, January 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185573.

Kegley:1959:RPF

[643] Charles W. Kegley. Reflections on Philipp Frank’s philosophy of science. Philosophy of Science, 26(1):35–40, January 1959. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185574.

Schlesinger:1959:PSV

[644] G. Schlesinger. The principle of simplicity and verifiability. Philosophy of Science, 26(1):41–42, January 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185575.

Feuer:1959:RPS

[645] Lewis S. Feuer. Rejoinder on the principle of simplicity. Philosophy of Science, 26(1):43–45, January 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185576.

Dunham:1959:BRBa

[646] Barrows Dunham. Book review: The Origins of Marxian Thought by Auguste Cornu. Philosophy of Science, 26(1):46–47, January 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185577.

Gluck:1959:BRBa

[647] Samuel E. Gluck. Book reviews: Nonparametric Methods in Statistics by D. A. S. Fraser; Nonparametric Statistics for the Behavioral Sciences by Sidney Siegel. Philosophy of Science, 26(1):47–48, January 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185578.

Gluck:1959:BRBb

[648] Samuel E. Gluck. Book review: Logical Design of Digital Computers by Montgomery Phister. Philosophy of Science, 26(1):48–50, January 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185579.

Parsons:1959:BRB

[649] Howard L. Parsons. Book review: The Promise of Modern Life; An Interrelational View by D. W. Gotshalk. Philosophy of Science, 26(1):50–51, January 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185580.

Frankfurt:1959:BRB

[650] Harry G. Frankfurt. Book review: On Shame and the Search for Identity by Helen Merrell Lynd. Philosophy of Science, 26(1):51–52, January 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185581.

Burgener:1959:BRBa

[651] R. J. C. Burgener. Book review: The Anatomy of the World by Michael Macklem. Philosophy of Science, 26(1):52–53, January 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185582.

Haworth:1959:BRB

[652] Lawrence Haworth. Book review: Social Theory and Social Structure by Robert Merton. Philosophy of Science, 26(1):53, January 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185583.

Lucca:1959:BRBa

[653] John De Lucca. Book review: In Search of Man by Andre Missenard; Lawrence G. Blochman. Philosophy of Science, 26(1):53–54, January 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185584.

Lucca:1959:BRBb

[654] John De Lucca. Book review: In Search of Reality by Viscount Samuel. Philosophy of Science, 26(1):54–55, January 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.

jstor.org/stable/185585.

Katz:1959:BRB

[655] Joseph Katz. Book review: Culture, Psychiatry and Human Values; The Methods and Values of a Social Psychiatry by Marvin K. Opler. Philosophy of Science, 26(1):55–57, January 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185586.

Grossmann:1959:BRB

[656] Reinhardt Grossmann. Book review: Semantics and Necessary Turth by Arthur Pap. Philosophy of Science, 26(1):57–58, January 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185587.

Fy:1959:BRB

[657] Curt Fy. Book review: Experimental Psychology and Other Essays by Ivan Petrovich Pavlov. Philosophy of Science, 26(1):58–59, January 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185588.

Anonymous:1959:BR

[658] Anonymous. Books received. Philosophy of Science, 26(1):60–61, January 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185589.

Anonymous:1959:ADa

[659] Anonymous. Abstracts from “dialectica”. Philosophy of Science, 26(1): 62, January 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185590.

Anonymous:1959:BMa

[660] Anonymous. Back matter. Philosophy of Science, 26(1):63–76, January 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185591.

Anonymous:1959:FMa

[661] Anonymous. Front matter. Philosophy of Science, 26(1):??, January 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185569.

Cowan:1959:EE

[662] Thomas A. Cowan. Experience and experiment. Philosophy of Science, 26 (2):77–83, April 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185253.

Blalock:1959:TCS

[663] H. M. Blalock, Jr. and Ann B. Blalock. Toward a clarification of system analysis in the social sciences. Philosophy of Science, 26(2):84–92, April 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185254.

Hochberg:1959:PBP

[664] Herbert Hochberg. Physicalism, behaviorism and phenomena. Philosophy of Science, 26(2):93–103, April 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185255.

Herburt:1959:ASD

[665] George Krzywicki Herburt. The analytic and the synthetic. The Duhemian argument and some contemporary philosophers. Philosophy of Science, 26(2):104–113, April 1959. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185256.

Raab:1959:HFR

[666] Francis V. Raab. History, freedom, and responsibility. Philosophy of Science, 26(2):114–124, April 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185257.

Moon:1959:MP

[667] Parry Moon and Domina Eberle Spencer. Mach’s Principle. Philosophy of Science, 26(2):125–134, April 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185258.

Builder:1959:RCP

[668] Geoffrey Builder. The resolution of the clock paradox. Philosophy of Science, 26(2):135–144, April 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185259.

Rosen:1959:CPT

[669] Philip Rosen. The clock paradox and thermodynamics. Philosophy of Science, 26(2):145–147, April 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185260.

Copi:1959:BRB

[670] Irving M. Copi. Book review: Intention by G. E. M. Anscombe. Philosophy of Science, 26(2):148–149, April 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185261.

Hardin:1959:BRBa

[671] Clyde L. Hardin. Book reviews: A Modern Introduction to Logic by John W. Blyth; Principles of Right Reason by Henry S. Leonard. Philosophy of Science, 26(2):149–150, April 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185262.

Lande:1959:BRB

[672] Alfred Landé. Book review: Atomic Physics and Human Knowledge by Niels Bohr. Philosophy of Science, 26(2):150–153, April 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185263.

Johnstone:1959:BRBa

[673] Henry W. Johnstone, Jr. Book review: Concepts of Force: A Study in the Foundations of Dynamics by Max Jammer. Philosophy of Science, 26(2):153–155, April 1959. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185264.

Ore:1959:BRB

[674] Oystein Ore. Book review: The Exact Sciences in Antiquity by O. Neugebauer. Philosophy of Science, 26(2):155, April 1959. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185265.

Stannard:1959:BRBa

[675] Jerry Stannard. Book review: The Physical World of the Greeks by S. Sambursky; Merton Dagut. Philosophy of Science, 26(2):155–157, April 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185266.

Hinshaw:1959:BRB

[676] Virgil Hinshaw, Jr. Book review: The Sociology of Knowledge; An Essay in Aid of a Deeper Understanding of the History of Ideas by Werner Stark. Philosophy of Science, 26(2):157–160, April 1959. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185267.

Hardin:1959:BRBb

[677] Clyde Laurence Hardin. Book review: The Uses of Argument by Stephen E. Toulmin. Philosophy of Science, 26(2):160–163, April 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185268.

Newmark:1959:BRB

[678] Leonard Newmark. Book review: The Principles of Semantics by Stephen Ullman. Philosophy of Science, 26(2):163–164, April 1959. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185269.

Stannard:1959:BRBb

[679] Jerry Stannard. Book review: Science and Religion in Seventeenth-Century England by Richard S. Westfall. Philosophy of Science, 26(2): 164–165, April 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185270.

Salmon:1959:BRBa

[680] Wesley C. Salmon. Book review: The Logical Problem of Induction by Georg Henrik von Wright. Philosophy of Science, 26(2):166, April 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185271.

Goldstein:1959:BRB

[681] Leon J. Goldstein. Book review: Man and Culture; An Evaluation of the Work of Bronislaw Malinowski by Raymond Firth. Philosophy of Science, 26(2):167–169, April 1959. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185272.

Anonymous:1959:BMb

[682] Anonymous. Back matter. Philosophy of Science, 26(2):??, April 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185273.

Anonymous:1959:FMb

[683] Anonymous. Front matter. Philosophy of Science, 26(2):??, April 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185252.

Ducasse:1959:SPM

[684] C. J. Ducasse. A statement from the President to the members of the philosophy of science association. Philosophy of Science, 26(3):171, July 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/184946.

Leonard:1959:IIT

[685] Henry S. Leonard. Interrogatives, imperatives, truth, falsity and lies. Philosophy of Science, 26(3):172–186, July 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184947.

Bird:1959:PTM

[686] Otto Bird. Peirce’s theory of methodology. Philosophy of Science, 26(3): 187–200, July 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/184948.

Caws:1959:FDS

[687] Peter Caws. The functions of definition in science. Philosophy of Science, 26(3):201–228, July 1959. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184949.

Laucks:1959:WNR

[688] Irving F. Laucks. Was Newton right after all? Philosophy of Science, 26(3):229–239, July 1959. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184950.

Schlesinger:1959:TAM

[689] G. Schlesinger. Two approaches to mathematical and physical systems. Philosophy of Science, 26(3):240–250, July 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184951.

Workman:1959:ISQ

[690] Rollin W. Workman. Is indeterminism supported by quantum theory? Philosophy of Science, 26(3):251–259, July 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/184952.

Matheson:1959:AD

[691] Gordon Matheson. The antinomy of designation. Philosophy of Science, 26(3):260–269, July 1959. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184953.

Moore:1959:BRBa

[692] Edward C. Moore. Book review: Personal Knowledge: Towards a Post-Critical Philosophy by Michael Polanyi. Philosophy of Science, 26(3):270– 272, July 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184954.

Burgener:1959:BRBb

[693] R. J. C. Burgener. Book review: Descartes’s Rules for the Direction of the Mind by Harold H. Joachim; E. E. Harris. Philosophy of Science, 26(3):272–274, July 1959. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/184955.

Hartung:1959:BRB

[694] Frank E. Hartung. Book review: The Science of Culture: A Study of Man and Civilization by Leslie A. White. Philosophy of Science, 26(3): 274, July 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184956.

Bar-Hillel:1959:BRBa

[695] Yehoshua Bar-Hillel. Book review: Language as Choice and Chance by G. Herdan. Philosophy of Science, 26(3):275, July 1959. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/184957.

Anonymous:1959:BMc

[696] Anonymous. Back matter. Philosophy of Science, 26(3):??, July 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184958.

Anonymous:1959:FMc

[697] Anonymous. Front matter. Philosophy of Science, 26(3):??, July 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/184945.

Leonard:1959:AP

[698] Henry S. Leonard. Authorship and purpose. Philosophy of Science, 26 (4):277–294, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185363.

Hartley:1959:MTE

[699] R. V. L. Hartley. A mechanistic theory of extra-atomic physics. Philosophy of Science, 26(4):295–309, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185364.

Hinshaw:1959:DVC

[700] Virgil Hinshaw, Jr. Determinism versus continuity. Philosophy of Science, 26(4):310–324, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185365.

Hanson:1959:FCC

[701] Norwood Russell Hanson. Five cautions for the Copenhagen interpretation’s critics. Philosophy of Science, 26(4):325–337, October 1959. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185366.

Gregg:1959:DWP

[702] John R. Gregg. On deciding whether protistans are cells. Philosophy of Science, 26(4):338–346, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185367.

Dettering:1959:LSS

[703] Richard W. Dettering. Linguistic superfluity in science. Philosophy of Science, 26(4):347–354, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185368.

Jack:1959:RBC

[704] Henry Jack. Reply to Barker’s criticism of formalism. Philosophy of Science, 26(4):355–361, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185369.

Linhart:1959:DBC

[705] H. Linhart. Darlington’s “On the Confirmation of Laws”. Philosophy of Science, 26(4):362, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/ 185370. See [640].

Darlington:1959:RLa

[706] Jared L. Darlington. Reply to Linhart. Philosophy of Science, 26(4):363, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185371.

Leblanc:1959:PDC

[707] Hugues Leblanc. Professor Darlington and the confirmation of laws. Philosophy of Science, 26(4):364–366, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/

stable/185372. See [640].

Darlington:1959:RLb

[708] Jared L. Darlington. Reply to Leblanc. Philosophy of Science, 26(4):367– 368, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185373.

Anonymous:1959:ADb

[709] Anonymous. Abstracts from “dialectica”. Philosophy of Science, 26(4): 369–371, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/185374.

McClellan:1959:BRB

[710] James E. McClellan. Book review: Cultural Foundations of Education: An Interdisciplinary Approach by Theodore Brameld. Philosophy of Science, 26(4):372, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185375.

Kaelin:1959:BRB

[711] E. F. Kaelin. Book review: Science and the Idea of God by C. A. Coulson. Philosophy of Science, 26(4):372–373, October 1959. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185376.

Werkmeister:1959:BRBa

[712] W. H. Werkmeister. Book reviews: Philosophy of Science, the Link between Science and Philosophy by Philipp Frank; Physical Science and Physical Reality by Louis O. Kattsoff. Philosophy of Science, 26(4):374– 375, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185377.

Doney:1959:BRB

[713] Willis Doney. Book review: Three Theories of Emotion; Some Views on Philosophical Method by Erik Götlind. Philosophy of Science, 26(4):375– 376, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185378.

Johnstone:1959:BRBb

[714] Henry W. Johnstone, Jr. Book review: The Physicist’s Conception of Nature by Werner Heisenberg; Arnold J. Pomerans. Philosophy of Science, 26(4):376, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185379.

MacKinnon:1959:BRB

[715] William J. MacKinnon. Book review: The True Believer; Thoughts on the Nature of Mass Movements by Eric Hoffer. Philosophy of Science, 26 (4):376–378, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185380.

Maziarz:1959:BRB

[716] Edward A. Maziarz. Book review: The History of Mathematics by Joseph E. Hofmann; Frank Gaynor; Henrietta P. Midonick. Philosophy of Science, 26(4):378–379, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/185381.

Dunham:1959:BRBb

[717] Barrows Dunham. Book review: The Academic Mind: Social Scientists in a Time of Crisis by Paul F. Lazarsfeld; Wagner Thielens. Philosophy of Science, 26(4):379–381, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.org/stable/

185382.

Bar-Hillel:1959:BRBb

[718] Yehoshua Bar-Hillel. Book review: Truth and Denotation, a Study in Semantical Theory by Richard M. Martin. Philosophy of Science, 26(4): 381, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185383.

Werkmeister:1959:BRBb

[719] W. H. Werkmeister. Book reviews: Ethics by P. H. Nowell-Smith; Value and Obligation: The Foundations of an Empiricist Ethical Theory by Alexander Sesonske. Philosophy of Science, 26(4):382–385, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185384.

Moore:1959:BRBb

[720] Edward C. Moore. Book review: Disease, Life, and Man by Rudolf Virchow; Lelland J. Rather. Philosophy of Science, 26(4):386, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185385.

Salmon:1959:BRBb

[721] Wesley C. Salmon. Book review: Probability, Statistics and Truth by Richard von Mises; Hilda Geiringer. Philosophy of Science, 26(4):387– 388, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185386.

Stannard:1959:BRBc

[722] Jerry Stannard. Book review: The Philosophical Writings of Chauncey Wright, Representative Selections by Chauncey Wright; Edward H. Madden. Philosophy of Science, 26(4):388, October 1959. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185387.

Anonymous:1959:VI

[723] Anonymous. Volume information. Philosophy of Science, 26(4):389–393, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185361.

Anonymous:1959:BMd

[724] Anonymous. Back matter. Philosophy of Science, 26(4):394, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL http://www.jstor.org/stable/185388.

Anonymous:1959:FMd

[725] Anonymous. Front matter. Philosophy of Science, 26(4):??, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

http://www.jstor.org/stable/185362.